### v1

# A Characterization for Tightness of the Sparse Moment-SOS Hierarchy

Abstract.
This paper studies the sparse Moment-SOS hierarchy of relaxations for solving sparse polynomial optimization problems. We show that this sparse hierarchy is tight if and only if the objective can be written as a sum of sparse nonnegative polynomials, each of which belongs to the sum of the ideal and quadratic module generated by the corresponding sparse constraints. Based on this characterization, we give several sufficient conditions for the sparse Moment-SOS hierarchy to be tight. In particular, we show that this sparse hierarchy is tight under some assumptions such as convexity, optimality conditions or finiteness of constraining sets.
Key words and phrases:
polynomial optimization, sparsity, moment, sum of squares, tight relaxation
2020 Mathematics Subject Classification:
90C23, 65K10, 90C22
1. Introduction
Let $x\coloneqq(x_{1},\ldots,x_{n})$ be an $n$-dimensional vector variable in $\mathbb{R}^{n}$. Suppose $\Delta_{1},\ldots,\Delta_{m}$ are subsets of $[n]\coloneqq\{1,\ldots,n\}$. For $\Delta_{i}=\{j_{1},\ldots,j_{n_{i}}\}$, denote the subvector $x_{\Delta_{i}}\coloneqq(x_{j_{1}},\ldots,x_{j_{n_{i}}}).$ We consider the sparse polynomial optimization problem
(1.1) $\left\{\begin{array}[]{cl}\displaystyle\min_{x\in\mathbb{R}^{n}}&f(x)\coloneqq f_{1}(x_{\Delta_{1}})+\dots+f_{m}(x_{\Delta_{m}})\\ \mbox{s.t.}&h_{i}(x_{\Delta_{i}})=0,\,g_{i}(x_{\Delta_{i}})\geq 0,\,i=1,\ldots,m.\\ \end{array}\right.$
In the above, each $f_{i}$ is a polynomial and $h_{i},g_{i}$ are vectors of polynomials in $x_{\Delta_{i}}$. Throughout the paper, a minimizer for (1.1) means it is a global minimizer. We denote by $f_{\min}$ the minimum value of (1.1) and denote
(1.2) $K_{\Delta_{i}}\coloneqq\{x_{\Delta_{i}}\in\mathbb{R}^{n_{i}}:h_{i}(x_{\Delta_{i}})=0,g_{i}(x_{\Delta_{i}})\geq 0\}.$
The feasible set $K$ of (1.1) is
$K=\bigcap\limits_{i=1}^{m}\{x\in\mathbb{R}^{n}:x_{\Delta_{i}}\in K_{\Delta_{i}}\}.$
General polynomial optimization problems can be solved by the Moment-SOS hierarchy of semidefinite programming relaxations proposed by Lasserre [14]. It produces a sequence of lower bounds for $f_{\min}$, which converges to $f_{\min}$ under the archimedeanness. Moreover, under some classical sufficient optimality conditions, the Moment-SOS hierarchy in [14] is tight, i.e., it has finite convergence [29]. Throughout the paper, the method in [14] is called the dense Moment-SOS hierarchy. We refer to the books [5, 17, 19, 30] about this topic.
The Moment-SOS hierarchy has strong performance for solving polynomial optimization. However, a concern in its computational practice is that the sizes of the resulting semidefinite programs grow quickly as the relaxation order increases. To improve computational efficiency, it is important to exploit sparsity. In some literature, the sparsity pattern in (1.1) is called correlative sparsity [15, 33], to be distinguished from term sparsity [35, 36, 37]. Some sparsity patterns can be given by arithmetic-geometric mean inequalities [10, 24, 25]. In this paper, we focus on correlative sparsity.
For a relaxation order $k$, we consider the sparse SOS relaxation
(1.3) $\left\{\begin{array}[]{cl}\max&\gamma\\ \mbox{s.t.}&f-\gamma\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}.\end{array}\right.$
The dual problem of (1.3) is the sparse moment relaxation
(1.4) $\left\{\begin{array}[]{cl}\min&\langle f,y\rangle\coloneqq\langle f_{1},y_{\Delta_{1}}\rangle+\cdots+\langle f_{m},y_{\Delta_{m}}\rangle\\ \mbox{s.t.}&\mathscr{V}_{h_{i}}^{spa,2k}[y]=0,\,L_{g_{i}}^{spa,k}[y]\succeq 0\ (i\in[m]),\\ &M_{spa}^{(k)}[y]\succeq 0,\,y_{0}=1,\,y\in\mathbb{R}^{\mathbb{U}_{k}}.\end{array}\right.$
We refer to Section 2 for the meaning of above notation. The sequence of relaxations (1.3)-(1.4) is called the sparse Moment-SOS hierarchy. The sparse relaxation has positive semidefinite (psd) matrix constraints whose sizes are much smaller than those of the dense relaxation. The sets $\Delta_{1},\ldots,\Delta_{m}$ are said to satisfy the running intersection property (RIP) if for every $j=2,\ldots,m$, it holds
(1.5) $\Delta_{j}\cap(\Delta_{1}\cup\cdots\cup\Delta_{j-1})\subseteq\Delta_{t}$
for some $t\leq j-1$. Let $f_{k}^{spa}$ and $f_{k}^{smo}$ denote the minimum values of (1.3) and (1.4) respectively. When RIP holds and each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}]$ is archimedean, the sequence of $f_{k}^{spa}$ converges to the minimum value $f_{\min}$ of (1.1) asymptotically [15]. A convergence rate of the sparse Moment-SOS hierarchy is given in the recent work [13]. We refer to [8, 12, 15, 22, 31, 32, 33] for related work on the sparse polynomial optimization. Recently, sparsity has also been exploited to solve noncommutative polynomial optimization [11, 34]. Some applications can be found in [26, 38, 39, 40, 41]. Sparse Moment-SOS relaxations can be implemented in the software TSSOS [21, 35, 37].
In practice, people often observe that the sparse Moment-SOS hierarchy has finite convergence, i.e., it is tight. However, there exist examples for which the dense hierarchy is tight while the sparse one is not (see [31]). To the best of the authors’ knowledge, there is very few work to characterize when the sparse Moment-SOS hierarchy is tight.
Contributions
This paper characterizes tightness of the sparse Moment-SOS hierarchy (1.3)-(1.4). Our main contributions are:
  • We give a sufficient and necessary condition for the sparse Moment-SOS hierarchy to be tight. More precisely, when the optimal value of (1.3) is achievable, we show that $f_{k}^{spa}=f_{\min}$ if and only if there exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ such that
    (1.6) $\boxed{\begin{gathered}p_{1}+\cdots+p_{m}+f_{\min}=0,\\ f_{i}+p_{i}\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2k}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2k},\ i\in[m].\end{gathered}}$
    We remark that the first equation in (1.6) is equivalent to
    $f-f_{\min}\,=\,(f_{1}+p_{1})+\cdots+(f_{m}+p_{m}).$
  • Under certain conditions, we show that the tightness of sparse Moment-SOS relaxations can be certified when flat truncations hold for the minimizer of the sparse moment relaxation (1.4). This also gives minimizers for (1.1).
  • For convex sparse polynomial optimization problems, we show that the sparse Moment-SOS hierarchy is tight under some general conditions. In particular, we show that the sparse relaxations are tight for all relaxation orders, when (1.1) is an SOS-convex optimization problem.
  • Based on the characterization, we prove the sparse Moment-SOS hierarchy is tight when some classical sufficiency optimality conditions hold or when each individual equality constraining variety is finite. In particular, when the RIP holds, we show that this sparse hierarchy is tight if each individual constraining variety is finite.
  • Based on the characterization, we prove the Schmüdgen type sparse Moment-SOS hierarchy is tight under some assumptions. In particular, we show that this sparse hierarchy is tight when the RIP holds and each individual constraining set is finite.
This paper is organized as follows. Some basics on polynomial optimization and algebraic geometry are reviewed in Section 2. In Section 3, we give a characterization for the tightness of sparse Moment-SOS hierarchy, and we study how to certify the tightness and get minimizers. Section 4 gives sufficient conditions for the sparse Moment-SOS hierarchy to be tight. The tightness of Schmüdgen type sparse Moment-SOS relaxations is investigated in Section 5. Some numerical experiments are presented in Section 6. Section 7 draws conclusions and makes some discussions.
2. Preliminaries
Notation
Denote by $\mathbb{N}$ the set of nonnegative integers and $\mathbb{R}$ the real field. For a positive integer $k$, let $[k]\coloneqq\{1,\ldots,k\}$. For a subset $\Delta_{i}\subseteq[n]$, denote by $\mathbb{R}^{\Delta_{i}}$ the space of real vectors $x_{\Delta_{i}}$. Let $\mathfrak{p}_{i}$ denote the projection from $\mathbb{R}^{n}$ to $\mathbb{R}^{\Delta_{i}}$ such that $\mathfrak{p}_{i}(x)=x_{\Delta_{i}}$ for all $x\in\mathbb{R}^{n}$. For $i,j\in[n]$, denote $\Delta_{ij}\coloneqq\Delta_{i}\cap\Delta_{j}$ and the projection $\mathfrak{p}_{ij}:\mathbb{R}^{\Delta_{i}}\rightarrow\mathbb{R}^{\Delta_{ij}}$ be such that $\mathfrak{p}_{ij}(x_{\Delta_{i}})=x_{\Delta_{ij}}$. The ring of polynomials in $x=(x_{1},\ldots,x_{n})$ with real coefficients is denoted as $\mathbb{R}[x]$. For a scalar or vector $p$ of polynomials, $\deg(p)$ denotes the maximal degree of its terms. For a power $\alpha=(\alpha_{1},\ldots,\alpha_{n})$, denote the monomial $x^{\alpha}\coloneqq x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$. For a degree $d$, the subset of polynomials in $\mathbb{R}[x]$ with degrees at most $d$ is denoted as $\mathbb{R}[x]_{d}$. The notation $\mathbb{R}[x_{\Delta_{i}}]$ and $\mathbb{R}[x_{\Delta_{i}}]_{d}$ are defined similarly. For $\phi\in\mathbb{R}[x]$ and subsets $A,B\subseteq\mathbb{R}[x]$, denote
$\phi A\,\coloneqq\,\{\phi p:p\in A\},\quad A+B\,\coloneqq\,\{p_{1}+p_{2}:p_{1}\in A,\,p_{2}\in B\}.$
For $f\in\mathbb{R}[x]$, $\nabla f$ denotes its gradient with respect to $x$, and $\nabla_{x_{\Delta_{i}}}f$ denotes the gradient with respect to $x_{\Delta_{i}}$. The Hessian matrices $\nabla^{2}f$ and $\nabla^{2}_{x_{\Delta_{i}}}f$ are defined similarly. The Euclidean norm of $x$ is $\|x\|\coloneqq(|x_{1}|^{2}+\cdots+|x_{n}|^{2})^{1/2}$. Denote by $e_{i}$ the canonical basis vector such that the $i$th entry is $1$ and 00 otherwise. Let $\mathcal{S}^{n}$ be the space of all $n\times n$ real symmetric matrices. For $X\in\mathcal{S}^{n}$, $X\succeq 0$ (resp., $X\succ 0$) means $X$ is positive semidefinite (resp., positive definite).
2.1. Ideals and SOS polynomials
A tuple $h_{i}\coloneqq(h_{i,1},\ldots,h_{i,\ell_{i}})$ of polynomials in $\mathbb{R}[x_{\Delta_{i}}]$ generates two ideals respectively in $\mathbb{R}[x_{\Delta_{i}}]$ and $\mathbb{R}[x]$ as
$\begin{array}[]{rcl}\operatorname{Ideal}_{\Delta_{i}}[h_{i}]&\coloneqq&h_{i,1}\mathbb{R}[x_{\Delta_{i}}]+\cdots+h_{i,\ell_{i}}\mathbb{R}[x_{\Delta_{i}}],\\ \operatorname{Ideal}[h_{i}]&\coloneqq&h_{i,1}\mathbb{R}[x]+\cdots+h_{i,\ell_{i}}\mathbb{R}[x].\end{array}$
The degree-$2k$ truncation of $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]$ is
$\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2k}\coloneqq h_{i,1}\mathbb{R}[x_{\Delta_{i}}]_{2k-\deg(h_{i,1})}+\cdots+h_{i,\ell_{i}}\mathbb{R}[x_{\Delta_{i}}]_{2k-\deg(h_{i,\ell_{i}})}.$
The truncation $\operatorname{Ideal}[h_{i}]_{2k}$ is similarly defined. For $h\coloneqq(h_{1},\ldots,h_{m})$, with each $h_{i}$ a tuple of polynomials in $\mathbb{R}[x_{\Delta_{i}}]$, we denote
(2.1) $\left\{\begin{array}[]{rcl}\operatorname{Ideal}[h]_{spa}&\coloneqq&\operatorname{Ideal}_{\Delta_{1}}[h_{1}]+\cdots+\operatorname{Ideal}_{\Delta_{m}}[h_{m}],\\ \operatorname{Ideal}[h]_{spa,2k}&\coloneqq&\operatorname{Ideal}_{\Delta_{1}}[h_{1}]_{2k}+\cdots+\operatorname{Ideal}_{\Delta_{m}}[h_{m}]_{2k}.\end{array}\right.$
For a set $P\subseteq\mathbb{R}[x_{\Delta_{i}}]$, its real variety is
$V_{\mathbb{R}}(P)\coloneqq\{x_{\Delta_{i}}\in\mathbb{R}^{\Delta_{i}}:p(x_{\Delta_{i}})=0\,\,\forall p\in P\}.$
The vanishing ideal of a subset $V_{i}\subseteq\mathbb{R}^{\Delta{i}}$ is
$I(V_{i})\coloneqq\{p\in\mathbb{R}[x_{\Delta_{i}}]:p(x_{\Delta_{i}})=0\,\,\forall x_{\Delta_{i}}\in V_{i}\}.$
Clearly, $P\subseteq I(V_{\mathbb{R}}(P))$. An ideal $J$ is said to be real radical if $J=I(V_{\mathbb{R}}(J))$.
A polynomial $\sigma\in\mathbb{R}[x_{\Delta_{i}}]$ is said to be a sum of squares (SOS) if there exist polynomials $p_{1},\ldots,p_{s}\in\mathbb{R}[x_{\Delta_{i}}]$ such that $\sigma=p_{1}^{2}+\cdots+p_{s}^{2}$. The cone of all SOS polynomials in $\mathbb{R}[x_{\Delta_{i}}]$ is denoted as $\Sigma[x_{\Delta_{i}}]$, and its degree-$2k$ truncation is
$\Sigma[x_{\Delta_{i}}]_{2k}\coloneqq\Sigma[x_{\Delta_{i}}]\cap\mathbb{R}[x_{\Delta_{i}}]_{2k}.$
The notation $\Sigma[x]$ and $\Sigma[x]_{2k}$ are similarly defined. A tuple $g_{i}\coloneqq(g_{i,1},\ldots,g_{i,s_{i}})$ of polynomials in $\mathbb{R}[x_{\Delta_{i}}]$ generates the quadratic modules in $\mathbb{R}[x_{\Delta_{i}}]$ and $\mathbb{R}[x]$ respectively as
$\begin{array}[]{rcl}\operatorname{QM}_{\Delta_{i}}[g_{i}]&\coloneqq&\Sigma[x_{\Delta_{i}}]+g_{i,1}\Sigma[x_{\Delta_{i}}]+\cdots+g_{i,s_{i}}\Sigma[x_{\Delta_{i}}],\\ \operatorname{QM}[g_{i}]&\coloneqq&\Sigma[x]+g_{i,1}\Sigma[x]+\cdots+g_{i,s_{i}}\Sigma[x].\end{array}$
The preorderings of $g_{i}$ in $\mathbb{R}[x_{\Delta_{i}}]$ and $\mathbb{R}[x]$ are respectively
$\operatorname{Pre}_{\Delta_{i}}[g_{i}]\,\coloneqq\sum_{J\subseteq[s_{i}]}\prod_{j\in J}g_{i,j}\Sigma[x_{\Delta_{i}}],\quad\operatorname{Pre}[g_{i}]\,\coloneqq\sum_{J\subseteq[s_{i}]}\prod_{j\in J}g_{i,j}\Sigma[x].$
The product in the above is $1$ if $J$ is empty. For an even degree $2k$, we denote the truncation ($g_{i,0}=1$)
$\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2k}\coloneqq\Big{\{}\sum_{j=0}^{s_{i}}\sigma_{j}g_{i,j}:\sigma_{j}\in\Sigma[x_{\Delta_{i}}],\deg(\sigma_{j}g_{i,j})\leq 2k\Big{\}}.$
The truncation $\operatorname{Pre}_{\Delta_{i}}[g_{i}]_{2k}$ is defined similarly. For $g\coloneqq(g_{1},\ldots,g_{m})$ with each $g_{i}$ a tuple of polynomials in $\mathbb{R}[x_{\Delta_{i}}]$, we denote
(2.2) $\left\{\begin{array}[]{rcl}\operatorname{QM}[g]_{spa}&\coloneqq&\operatorname{QM}_{\Delta_{1}}[g_{1}]+\cdots+\operatorname{QM}_{\Delta_{m}}[g_{m}],\\ \operatorname{Pre}[g]_{spa}&\coloneqq&\operatorname{Pre}_{\Delta_{1}}[g_{1}]+\ldots+\operatorname{Pre}_{\Delta_{m}}[g_{m}],\\ \operatorname{QM}[g]_{spa,2k}&\coloneqq&\operatorname{QM}_{\Delta_{1}}[g_{1}]_{2k}+\cdots+\operatorname{QM}_{\Delta_{m}}[g_{m}]_{2k},\\ \operatorname{Pre}[g]_{spa,2k}&\coloneqq&\operatorname{Pre}_{\Delta_{1}}[g_{1}]_{2k}+\ldots+\operatorname{Pre}_{\Delta_{m}}[g_{m}]_{2k}.\end{array}\right.$
The set $\operatorname{Ideal}_{\Delta_{i}}{[h_{i}]}+\operatorname{QM}_{\Delta_{i}}{[g_{i}]}$ is said to be archimedean if it contains a polynomial $q$ such that the set of all points $x_{\Delta_{i}}$ satisfying $q(x_{\Delta_{i}})\geq 0$ is compact. When $\operatorname{Ideal}_{\Delta_{i}}{[h_{i}]}+\operatorname{QM}_{\Delta_{i}}{[g_{i}]}$ is archimedean, the set $K_{\Delta_{i}}$ must be compact. The converse is not necessarily true. However, if $K_{\Delta_{i}}$ is compact, i.e., there exists $R>0$ such that $\|x_{\Delta_{i}}\|^{2}\leq R$ for all $x_{\Delta_{i}}\in K_{\Delta_{i}}$, then $\operatorname{Ideal}_{\Delta_{i}}{[h_{i}]}+\operatorname{QM}_{\Delta_{i}}[g_{i},R-\|x_{\Delta_{i}}\|^{2}]$ is archimedean. The following lemma is useful.
Lemma 2.1.
Let $f=f_{1}+\dots+f_{m}$ be such that each $f_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ and let $G=G_{1}+\cdots+G_{m}$ be such that each $G_{i}\subseteq\mathbb{R}[x_{\Delta_{i}}]$. Then, for a given $\gamma\in\mathbb{R}$, it holds $f-\gamma\in G$ if and only if there exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ such that
(2.3) $\boxed{\begin{gathered}p_{1}+\cdots+p_{m}+\gamma=0,\\ f_{i}+p_{i}\in G_{i},\,i=1,\ldots,m.\end{gathered}}$
Proof.
The “if” implication is straightforward. For the “only if” implication, suppose $f-\gamma\in G,$ then there exist $s_{i}\in G_{i}$ satisfying
(2.4) $f-\gamma=s_{1}+\cdots+s_{m}.$
For each $i$, let $p_{i}\,\coloneqq\,s_{i}-f_{i}.$ Since $f_{i},s_{i}\in\mathbb{R}[x_{\Delta_{i}}]$, we have $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ and
$p_{1}+\cdots+p_{m}$ $=s_{1}+\cdots+s_{m}-(f_{1}+\cdots+f_{m})$
$=s_{1}+\cdots+s_{m}-f=-\gamma.$
Therefore, $p_{1}+\cdots+p_{m}+\gamma=0$ and $f_{i}+p_{i}=s_{i}\in G_{i}$ for all $i$. So, (2.3) holds. ∎
The feasible set of the sparse optimization problem (1.1) is
(2.5) $K=\left\{x\in\mathbb{R}^{n}\left|\begin{array}[]{c}h_{i}(x_{\Delta_{i}})=0,\ g_{i}(x_{\Delta_{i}})\geq 0,\\ i=1,\ldots,m\end{array}\right.\right\}.$
In some applications, one is interested in a certificate for $K$ to be empty. The Positivstellensatz (see [2]) for $K=\emptyset$ is
(2.6) $1+\sigma+\phi=0,$
where $\sigma\in\operatorname{Pre}[g]\coloneqq\sum_{i=1}^{m}\operatorname{Pre}[g_{i}]$ and $\phi\in\operatorname{Ideal}[h]\coloneqq\sum_{i=1}^{m}\operatorname{Ideal}[h_{i}]$. The Positivstellensatz (2.6) is said to be sparse if (2.6) holds for $\sigma\in\operatorname{Pre}[g]_{spa}$ and $\phi\in\operatorname{Ideal}[h]_{spa}$.
Theorem 2.2.
Let $K$ be the set as in (2.5). Then the sparse Positivstellensatz (2.6) holds if and only if there exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ such that
(2.7) $\boxed{\begin{array}[]{c}p_{1}+\cdots+p_{m}=0,\\ -1+p_{i}\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{Pre}_{\Delta_{i}}[g_{i}],\,\,i\in[m].\end{array}}$
Proof.
The conclusion follows from Lemma 2.1 with $\gamma=0$, $f_{i}=-1$ and $G_{i}=\operatorname{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{Pre}_{\Delta_{i}}[g_{i}]$. ∎
2.2. Sparse moments
Denote the set of monomial powers in $x_{\Delta_{i}}$ as
$\mathbb{N}^{\Delta_{i}}\coloneqq\{\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}^{n}:\alpha_{j}=0\,\,\forall j\notin\Delta_{i}\}.$
For a degree $d$, denote $\mathbb{N}_{d}^{\Delta_{i}}\coloneqq\{\alpha\in\mathbb{N}^{\Delta_{i}}:|\alpha|\leq d\},$ where $|\alpha|\coloneqq\alpha_{1}+\cdots+\alpha_{n}$. The vector of all monomials in $x_{\Delta_{i}}$ in the graded lexicographic order with degrees up to $d$ is denoted as $[x_{\Delta_{i}}]_{d}$, i.e.,
(2.8) $[x_{\Delta_{i}}]_{d}\,=\,\big{(}x^{\alpha}\big{)}_{\alpha\in\mathbb{N}_{d}^{\Delta_{i}}}.$
Denote by $\mathbb{R}^{\mathbb{N}_{d}^{\Delta_{i}}}$ the space of all real vectors labeled by $\alpha\in\mathbb{N}_{d}^{\Delta_{i}}$. A vector $y_{\Delta_{i}}$ in $\mathbb{R}^{\mathbb{N}_{d}^{\Delta_{i}}}$ is called a truncated multi-sequence (tms) of degree $d$. For a given $y_{\Delta_{i}}\in\mathbb{R}^{\mathbb{N}_{d}^{\Delta_{i}}}$, the Riesz functional generated by $y_{\Delta_{i}}$ is the linear functional $\mathscr{R}_{y_{\Delta_{i}}}$ acting on $\mathbb{R}[x_{\Delta_{i}}]$ such that
$\mathscr{R}_{y_{\Delta_{i}}}(x^{\alpha})\,=\,(y_{\Delta_{i}})_{\alpha}\quad\mbox{for each}\quad\alpha\in\mathbb{N}_{d}^{\Delta_{i}}.$
This induces the bilinear operation
(2.9) $\langle p,y_{\Delta_{i}}\rangle\coloneqq\mathscr{R}_{y_{\Delta_{i}}}(p).$
The localizing vector and the localizing matrix of $p\in\mathbb{R}[x_{\Delta_{i}}]$, generated by $y_{\Delta_{i}}$, are respectively
$\begin{array}[]{rcl}\mathscr{V}_{p}^{\Delta_{i},2k}[y_{\Delta_{i}}]&\coloneqq&\mathscr{R}_{y_{\Delta_{i}}}\left(p(x_{\Delta_{i}})[x_{\Delta_{i}}]_{k_{1}}\right),\\ L_{p}^{\Delta_{i},k}[y_{\Delta_{i}}]&\coloneqq&\mathscr{R}_{y_{\Delta_{i}}}\left(p(x_{\Delta_{i}})[x_{\Delta_{i}}]_{k_{2}}[x_{\Delta_{i}}]^{T}_{k_{2}}\right).\end{array}$
In the above, the Riesz functional is applied entry-wise and
$k_{1}\coloneqq 2k-\deg(p),\quad k_{2}\coloneqq\lfloor k-\deg(p)/2\rfloor.$
In particular, when $p=1$ is the constant polynomial, we get the moment matrix
$M_{\Delta_{i}}^{(k)}[y_{\Delta_{i}}]\coloneqq L_{1}^{\Delta_{i},k}[y_{\Delta_{i}}].$
Recall that $h_{i}=(h_{i,1},\ldots,h_{i,\ell_{i}})$ and $g_{i}\coloneqq(g_{i,1},\ldots,g_{i,s_{i}})$ are tuples of polynomials in $\mathbb{R}[x_{\Delta_{i}}]$. We denote
$\begin{array}[]{rcl}\mathscr{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}]&\coloneqq&(\mathscr{V}_{h_{i,1}}^{\Delta_{i},2k}[y_{\Delta_{i}}],\ldots,\mathscr{V}_{h_{i,\ell_{i}}}^{\Delta_{i},2k}[y_{\Delta_{i}}]),\\ L_{g_{i}}^{\Delta_{i},k}[y_{\Delta_{i}}]&\coloneqq&\mbox{diag}\left(L_{g_{i,1}}^{\Delta_{i},k}[y_{\Delta_{i}}],\ldots,L_{g_{i,s_{i}}}^{\Delta_{i},k}[y_{\Delta_{i}}]\right).\end{array}$
For a given $k$, denote the monomial power set
(2.10) $\mathbb{U}_{k}\coloneqq\bigcup_{i\in[m]}\mathbb{N}_{2k}^{\Delta_{i}}.$
Let $\mathbb{R}^{\mathbb{U}_{k}}$ denote the space of real vectors $y=(y_{\alpha})_{\alpha\in\mathbb{U}_{k}}.$ For given $y\in\mathbb{R}^{\mathbb{U}_{k}}$, we denote the subvector
(2.11) $y_{\Delta_{i}}\,\coloneqq\,(y_{\alpha})_{\alpha\in\mathbb{N}_{2k}^{\Delta_{i}}}.$
For the objective $f$ in (1.1), we denote
(2.12) $\langle f,y\rangle\,\coloneqq\,\langle f_{1},y_{\Delta_{1}}\rangle+\cdots+\langle f_{m},y_{\Delta_{m}}\rangle.$
For a degree $t\leq k$, the $2t$-degree truncation of $y_{\Delta_{i}}$ is the subvector
(2.13) $y_{\Delta_{i}}|_{2t}\,\coloneqq\,(y_{\alpha})_{\alpha\in\mathbb{N}_{2t}^{\Delta_{i}}}.$
The sparse localizing vectors/matrices are denoted as
(2.14) $\left\{\begin{array}[]{rcl}\mathscr{V}_{h_{i}}^{spa,2k}[y]&\coloneqq&\mathscr{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}],\\ L_{g_{i}}^{spa,k}[y]&\coloneqq&L_{g_{i}}^{\Delta_{i},k}[y_{\Delta_{i}}],\\ M^{(k)}_{spa}[y]&\coloneqq&\mbox{diag}\left(M_{\Delta_{1}}^{(k)}[y_{\Delta_{1}}],\ldots,M_{\Delta_{m}}^{(k)}[y_{\Delta_{m}}]\right).\end{array}\right.$
We refer to [5, 17, 19, 30] for more detailed introductions to polynomial optimization.
3. The characterization for tightness
In this section, we characterize when the sparse Moment-SOS hierarchy is tight for solving (1.1). Let
(3.1) $k_{0}\coloneqq\max_{i\in[m]}\left\{\lceil\deg(f)/2\rceil,\lceil\deg(g_{i})/2\rceil,\lceil\deg(h_{i})/2\rceil\right\}.$
For a degree $k\geq k_{0}$, the $k$th order sparse SOS relaxation for (1.1) is
(3.2) $\left\{\begin{array}[]{ccl}f_{k}^{spa}\coloneqq&\max&\gamma\\ &\mbox{s.t.}&f-\gamma\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}.\end{array}\right.$
Its dual optimization problem is the sparse moment relaxation
(3.3) $\left\{\begin{array}[]{rcl}f_{k}^{smo}\coloneqq&\min&\langle f,y\rangle=\langle f_{1},y_{\Delta_{1}}\rangle+\cdots+\langle f_{m},y_{\Delta_{m}}\rangle\\ &\mbox{s.t.}&\mathscr{V}_{h_{i}}^{spa,2k}[y]=0,L_{g_{i}}^{spa,k}[y]\succeq 0\ (i\in[m]),\\ &&M_{spa}^{(k)}[y]\succeq 0,y_{0}=1,\,y\in\mathbb{R}^{\mathbb{U}_{k}}.\end{array}\right.$
We refer to Subsections 2.1 and 2.2 for the above notation. Recall that $f_{\min}$ denotes the minimum value of (1.1). When the RIP holds, if each $\operatorname{Ideal}_{\Delta_{i}}{[h_{i}]}+\operatorname{QM}_{\Delta_{i}}{[g_{i}]}$ is archimedean, then $f_{k}^{spa}\rightarrow f_{\min}$ as $k\rightarrow\infty$, as shown in [15]. When $f_{k}^{spa}=f_{\min}$ for some $k$, the hierarchy of (3.2) is said to be tight. Similarly, if $f_{k}^{smo}=f_{\min}$ for some $k$, then the hierarchy of (3.3) is tight. If they are both tight, the sparse Moment-SOS hierarchy of (3.2)-(3.3) is said to be tight, or to have finite convergence.
3.1. Characterization for the tightness
First, we give a sufficient and necessary condition for the sparse Moment-SOS hierarchy of (3.2)-(3.3) to be tight. The following is the main theorem.
Theorem 3.1.
For the sparse Moment-SOS hierarchy of (3.2)-(3.3), we have:
  1. (i)
    For a relaxation order $k\geq k_{0}$, it holds
    (3.4) $f-f_{\min}\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}$
    if and only if there exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]_{2k}$ such that
    (3.5) $\boxed{\begin{gathered}p_{1}+\cdots+p_{m}+f_{\min}=0,\\ f_{i}+p_{i}\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2k}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2k},\ i\in[m].\end{gathered}}$
    The equation in the above is equivalent to that
    $f-f_{\min}\,=\,(f_{1}+p_{1})+\cdots+(f_{m}+p_{m}).$
  2. (ii)
    When (3.5) holds for some order $k$, the minimum value $f_{\min}$ of (1.1) is achievable if and only if all sparse polynomials $f_{i}+p_{i}$ have a common zero in $K$, i.e., there exists $u\in K$ such that $f_{i}(u_{\Delta_{i}})+p_{i}(u_{\Delta_{i}})=0$ for all $i\in[m]$.
Proof.
(i) Let $\gamma=f_{\min}$, $G=\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}$ and
$G_{i}=\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2k}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2k},\,\,i=1,\ldots,m.$
Then, the conclusion follows from Lemma 2.1.
(ii) Assume (3.5) holds for some order $k$.
($\Rightarrow$): Suppose $f_{\min}$ is achievable for (1.1), then there exists a minimizer $u\in K$ such that $f_{\min}=f(u)$. Note
$f(x)-f_{\min}$ $=$ $-\Big{[}f_{\min}+\sum_{i=1}^{m}p_{i}(x_{\Delta_{i}})\Big{]}+\sum_{i=1}^{m}[f_{i}(x_{\Delta_{i}})+p_{i}(x_{\Delta_{i}})]$
$\in$ $\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}.$
Since $p_{1}+\dots+p_{m}+f_{\min}=0,$ we have
$\sum_{i=1}^{m}(f_{i}(u_{\Delta_{i}})+p_{i}(u_{\Delta_{i}}))=0.$
Since each $f_{i}(u_{\Delta_{i}})+p_{i}(u_{\Delta_{i}})\geq 0$ on $K_{\Delta_{i}}$, we have $f_{i}(u_{\Delta_{i}})+p_{i}(u_{\Delta_{i}})=0$ for all $i\in[m]$. Therefore, $u\in K$ is a common zero of all $f_{i}+p_{i}$.
($\Leftarrow$): Suppose $u\in K$ is a common zero of all $f_{i}+p_{i}$, then
$\sum\limits_{i=1}^{m}f_{i}(u_{\Delta_{i}})+p_{i}(u_{\Delta_{i}})=f(u)+p_{1}(u_{\Delta_{1}})+\cdots+p_{m}(u_{\Delta_{m}})=0.$
Since $p_{1}+\cdots+p_{m}+f_{\min}=0$, $f(u)=f_{\min}$, so $f_{\min}$ is achievable. ∎
Theorem 3.1 gives a sufficient and necessary condition for the membership (3.4), which implies the tightness $f^{spa}_{k}=f_{\min}$. When the optimal value of (3.2) is achievable, $f^{spa}_{k}=f_{\min}$ if and only if (3.4) holds. When (3.2) does not achieve its optimal value, we may not have (3.4), even if the relaxation (3.2) is tight. For such cases, we have the following characterization for the tightness $f^{spa}_{k}=f_{\min}$.
Theorem 3.2.
The $k$th order sparse SOS relaxation (3.2) is tight (i.e., $f_{\min}=f^{spa}_{k}$) if and only if for every $\epsilon>0$, there exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]_{2k}$ such that
(3.6) $\boxed{\begin{gathered}p_{1}+\cdots+p_{m}+f_{\min}=0,\\ f_{i}+p_{i}+\epsilon\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2k}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2k},\ i\in[m].\end{gathered}}$
Proof.
($\Leftarrow$): Suppose (3.6) holds for some polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]_{2k}$, then
$f-(f_{\min}-m\epsilon)$ $=$ $-\Big{[}f_{\min}+\sum_{i=1}^{m}p_{i}\Big{]}+\sum_{i=1}^{m}(f_{i}+p_{i}+\epsilon)$
$\in$ $\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}.$
This means that $\gamma=f_{\min}-m\epsilon$ is feasible for the $k$th order sparse SOS relaxation (3.2), so $f_{k}^{spa}\geq f_{\min}-m\epsilon$. Since $\epsilon>0$ is arbitrary, we get $f_{k}^{spa}\geq f_{\min}$. On the other hand, we always have $f_{k}^{spa}\leq f_{\min}$, so $f_{k}^{spa}=f_{\min}$.
($\Rightarrow$): For an arbitrary $\epsilon>0$, let
$f^{\epsilon}_{i}(x_{\Delta_{i}})\,\coloneqq\,f_{i}(x_{\Delta_{i}})+\epsilon.$
Suppose the $k$th order relaxation (3.2) is tight. Then, every $\gamma<f_{\min}=f_{k}^{spa}$ is feasible for (3.2). This is because for every $\epsilon>0$, there exists a feasible $\hat{\gamma}$ such that $\hat{\gamma}>f_{\min}-m\epsilon$. So, $f-\hat{\gamma}\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}$ and
$f-(f_{\min}-m\epsilon)=f-\hat{\gamma}+[\hat{\gamma}-(f_{\min}-m\epsilon)]\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k},$
since $\hat{\gamma}-(f_{\min}-m\epsilon)>0$. Thus, $\gamma=f_{\min}-m\epsilon$ is feasible for (3.2), and
$f(x)-(f_{\min}-m\epsilon)=\sum_{i=1}^{m}f^{\epsilon}_{i}-f_{\min}\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}.$
We apply Lemma 2.1 with $\gamma=f_{\min}$ and $f_{i}=f_{i}^{\epsilon}$, then there exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]_{2k}$ such that $p_{1}+\cdots+p_{m}+f_{\min}=0$ and for all $i$,
$f_{i}^{\epsilon}+p_{i}=f_{i}+p_{i}+\epsilon\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2k}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2k}.$
So, (3.6) holds. ∎
3.2. Detecting tightness and extracting minimizers
We now discuss how to detect tightness of sparse Moment-SOS relaxations (3.2)-(3.3). The tightness of (3.2) can be detected if there exists a feasible point $u$ such that $f_{k}^{spa}=f(u)$; similarly, the tightness of (3.3) can be detected by $f_{k}^{smo}=f(u)$. This issue is related to how to extract a minimizer for (1.1) from the sparse moment relaxations.
For each $i\in[m]$, denote the degrees
(3.7) $\begin{array}[]{rcl}d_{i}&\coloneqq&\max\{\lceil\deg(h_{i})/2\rceil,\lceil\deg(g_{i})/2\rceil\}.\end{array}$
Suppose $y^{*}$ is a minimizer of (3.3) for a relaxation order $k\geq k_{0}$. To extract a minimizer, we typically need to consider a representing measure for the subvector $y^{*}_{\Delta_{i}}$. Suppose there exists a degree $t\in[k_{0},k]$ such that
(3.8) $\boxed{\begin{gathered}{y^{*}_{\Delta_{i}}|_{2t}=\lambda_{i,1}[u^{(i,1)}]_{2t}+\cdots+\lambda_{i,r_{i}}[u^{(i,r_{i})}]_{2t},}\\ \lambda_{i,1}>0,\ldots,\lambda_{i,r_{i}}>0,\quad\lambda_{i,1}+\cdots+\lambda_{i,r_{i}}=1\end{gathered}}$
for some points $u^{(i,j)}\in K_{\Delta_{i}}$. We refer to (2.8), (2.11) and (2.13) for the above notation. Denote the set
(3.9) $\mathcal{X}_{\Delta_{i}}\coloneqq\{u^{(i,1)},\ldots,u^{(i,r_{i})}\}.$
Theorem 3.3.
Suppose $y^{*}$ is a minimizer of (3.3) and there exists $t\in[k_{0},k]$ such that (3.8) holds for all $i\in[m]$. Let $x^{*}=(x_{1}^{*},\,\cdots,\,x_{n}^{*})$ be a point such that each $x^{*}_{\Delta_{i}}\in\mathcal{X}_{\Delta_{i}}$. If $f_{t}^{smo}=f_{k}^{smo}$, then $x^{*}$ is a minimizer of (1.1).
Proof.
Since (3.8) holds, we have
$M_{\Delta_{i}}^{(t)}[y^{*}_{\Delta_{i}}]=\rho_{i}[x_{\Delta_{i}}^{*}]_{t}[x_{\Delta_{i}}^{*}]_{t}^{T}+W_{\Delta_{i}}$
for a positive scalar $\rho_{i}>0$ and a moment matrix $W_{\Delta_{i}}\succeq 0$. Let
$\rho\coloneqq\min_{1\leq i\leq m}\rho_{i}.$
Since $(y^{*}_{\Delta_{i}})_{0}=1$, we have $0<\rho\leq\rho_{i}\leq 1$ for every $i\in[m]$, hence
$\widehat{W}_{\Delta_{i}}\,\coloneqq\,(\rho_{i}-\rho)[x_{\Delta_{i}}^{*}]_{t}[x_{\Delta_{i}}^{*}]_{t}^{T}+W_{\Delta_{i}}\succeq 0.$
Note that $\widehat{W}_{\Delta_{i}}$ and $W_{\Delta_{i}}$ are also moment matrices for some tms’s.
For the case $\rho=1$, we have $\widehat{W}_{\Delta_{i}}=W_{\Delta_{i}}=0$ for all $i$. So, $\mathcal{X}_{\Delta_{i}}$ consists of the single point $x_{\Delta_{i}}^{*}$, and hence $y^{*}_{\Delta_{i}}|_{2t}=[x_{\Delta_{i}}^{*}]_{2t}$ for all $i\in[m]$. Then, one can see that
$f_{k}^{smo}=\langle f,y^{*}\rangle=\sum_{i=1}^{m}\langle f_{i},y^{*}_{\Delta_{i}}\rangle=\sum_{i=1}^{m}f_{i}(x_{\Delta_{i}}^{*})=f(x^{*}).$
Since each $x_{\Delta_{i}}^{*}\in K_{\Delta_{i}}$, we know $x^{*}$ is a minimizer of (1.1).
For the case $10<\rho<$ 1, define the new tms
$\hat{y}\coloneqq(\hat{y}_{\alpha})_{\alpha\in\mathbb{U}_{k}}\quad\text{where each}\quad\hat{y}_{\alpha}=\left(x_{\Delta_{i}}^{*}\right)^{\alpha}.$
Let $\tilde{y}\coloneqq(\tilde{y}_{\alpha})_{\alpha\in\mathbb{U}_{k}}$ be the tms such that $y^{*}=\rho\hat{y}+(1-\rho)\tilde{y}$, then
$\mathscr{V}_{h_{i}}^{spa,2t}[y^{*}]$ $=$ $\rho\mathscr{V}_{h_{i}}^{spa,2t}[\hat{y}]+(1-\rho)\mathscr{V}_{h_{i}}^{spa,2t}[\tilde{y}],$
$L_{g_{i}}^{spa,t}[y^{*}]$ $=$ $\rho L_{g_{i}}^{spa,t}[\hat{y}]+(1-\rho)L_{g_{i}}^{spa,t}[\tilde{y}],$
$M_{\Delta_{i}}^{(t)}[y^{*}_{\Delta_{i}}]$ $=$ $\rho M_{\Delta_{i}}^{(t)}[\hat{y}_{\Delta_{i}}]+(1-\rho)M_{\Delta_{i}}^{(t)}[\tilde{y}_{\Delta_{i}}].$
Both $\hat{y}$ and $\tilde{y}$ are feasible for (3.3) with the relaxation order equal to $t$, because
$\mathscr{V}_{h_{i}}^{spa,2t}[\tilde{y}]$ $=$ $\frac{1}{1-\rho}\left(\mathscr{V}_{h_{i}}^{spa,2t}[y^{*}]-\rho\mathscr{V}_{h_{i}}^{spa,2t}[\hat{y}]\right)=0,$
$L_{g_{i}}^{spa,t}[\tilde{y}]$ $=$ $\frac{1}{1-\rho}\left(L_{g_{i}}^{spa,t}[y^{*}]-\rho L_{g_{i}}^{spa,t}[\hat{y}]\right)\succeq 0,$
$M_{\Delta_{i}}^{(t)}[\tilde{y}_{\Delta_{i}}]$ $=$ $\frac{1}{1-\rho}\left(M_{\Delta_{i}}^{(t)}[y^{*}_{\Delta_{i}}]-\rho M_{\Delta_{i}}^{(t)}[\hat{y}_{\Delta_{i}}]\right)\succeq 0.$
Since $f_{t}^{smo}=f_{k}^{smo}$, the truncation $y^{*}|_{2t}$ is a minimizer for (3.3) with $k=t$, so
$\langle f,y^{*}\rangle\leq\langle f,\hat{y}\rangle,\quad\langle f,y^{*}\rangle\leq\langle f,\tilde{y}\rangle.$
On the other hand, it also holds that
$f_{t}^{smo}=\langle f,y^{*}\rangle=\rho\langle f,\hat{y}\rangle+(1-\rho)\langle f,\tilde{y}\rangle.$
Since $10<\rho<$ 1, we must have
$f_{t}^{smo}=\langle f,y^{*}\rangle=\langle f,\hat{y}\rangle=\langle f,\tilde{y}\rangle.$
Note that $f(x^{*})=\langle f,\hat{y}\rangle$, so $f(x^{*})=f_{t}^{smo}$. Since $f_{k}^{smo}\leq f_{\min}$, we have
$f(x^{*})=f_{t}^{smo}\leq f_{\min}\leq f(x^{*}).$
This shows that $x^{*}$ is a minimizer of (1.1). ∎
For the decomposition (3.8) to hold, we typically need to assume the flat truncation condition: there exists $t\in[k_{0},k]$ such that ($d_{i}$ is given in (3.7))
(3.10) $\mbox{rank}\,M_{\Delta_{i}}^{(t)}[y^{*}_{\Delta_{i}}]=\mbox{rank}\,M_{\Delta_{i}}^{(t-d_{i})}[y^{*}_{\Delta_{i}}].$
When (3.10) holds, we have the decomposition (3.8) with $r_{i}=\mbox{rank}\,M_{\Delta_{i}}^{(t)}[y^{*}_{\Delta_{i}}]$. We refer to [6, 18] for this fact. Under some genericity assumptions, the dense Moment-SOS hierarchy is tight if and only if the flat truncation holds (see [27, Theorem 2.2] or [30, Section 5.3]).
The connectedness of the sparsity pattern set $\mathcal{D}\coloneqq\{\Delta_{1},\ldots,\Delta_{m}\}$ can be given by the graph $G=([m],E)$ whose vertex set is $[m]$ and whose edge set $E$ consists of pairs $(i,j)$ such that $\Delta_{i}\cap\Delta_{j}\neq\emptyset.$ The set $\mathcal{D}$ is said to be a connected cover of $[n]$ if $G$ is a connected graph and $\Delta_{1}\cup\cdots\cup\Delta_{m}=[n]$. The following gives conditions for the existence of a representing measure for a truncation of $y^{*}$. For a degree $t\leq k$, denote (refer to (2.10) for the notation of $\mathbb{U}_{t}$)
(3.11) $y^{*}|_{spa,2t}=\big{(}y^{*}_{\alpha}\big{)}_{\alpha\in\mathbb{U}_{t}},\quad[u]_{spa,2t}=\big{(}u^{\alpha}\big{)}_{\alpha\in\mathbb{U}_{t}}.$
Theorem 3.4.
Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$. Assume $y^{*}$ is a minimizer of (3.3) and (3.10) holds for each $i$. Let $r_{i}$ be as in (3.8). Assume $r_{1}=\cdots=r_{m}=r$. If for all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$ it holds
(3.12) $r=\mbox{rank}\,M_{\Delta_{ij}}^{(t)}[y^{*}_{\Delta_{ij}}]=\mbox{rank}\,M_{\Delta_{ij}}^{(t-1)}[y^{*}_{\Delta_{ij}}],$
then there are points $u^{(1)},\ldots,u^{(r)}\in K$ such that
(3.13) $\boxed{\begin{gathered}y^{*}|_{spa,2t}=\lambda_{1}[u^{(1)}]_{spa,2t}+\cdots+\lambda_{r}[u^{(r)}]_{spa,2t},\\ \lambda_{1}>0,\ldots,\lambda_{r}>0,\,\lambda_{1}+\cdots+\lambda_{r}=1.\end{gathered}}$
Moreover, $u^{(1)},\ldots,u^{(r)}$ are minimizers of (1.1).
Proof.
Pick arbitrary $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, then (3.12) holds, by the given assumption. The flat truncation (3.10) implies the decomposition (3.8). Note that
$\begin{gathered}y^{*}_{\Delta_{ij}}|_{2t}=\lambda_{i,1}[\mathfrak{p}_{ij}(u^{(i,1)})]_{2t}+\cdots+\lambda_{i,r}[\mathfrak{p}_{ij}(u^{(i,r)})]_{2t},\\ y^{*}_{\Delta_{ji}}|_{2t}=\lambda_{j,1}[\mathfrak{p}_{ji}(u^{(j,1)})]_{2t}+\cdots+\lambda_{j,r}[\mathfrak{p}_{ji}(u^{(j,r)})]_{2t}.\end{gathered}$
Note $y^{*}_{\Delta_{ij}}|_{2t}=y^{*}_{\Delta_{ji}}|_{2t}$, since they are common entries of $y^{*}$. The condition (3.12) implies that $y^{*}_{\Delta_{ij}}|_{2t}$ and $y^{*}_{\Delta_{ji}}|_{2t}$ have the same unique representing measure, whose support consists of $r$ distinct points. This is shown in [3, 6, 18]. So, $\mathfrak{p}_{ij}(u^{(i,1)})$, …, $\mathfrak{p}_{ij}(u^{(i,r)})$ are distinct points in $\mathfrak{p}_{ij}(K_{\Delta_{i}})$, and $\mathfrak{p}_{ji}(u^{(j,1)})$, …, $\mathfrak{p}_{ji}(u^{(j,r)})$ are distinct points in $\mathfrak{p}_{ji}(K_{\Delta_{j}})$. Since the representing measure is unique,
$\{\mathfrak{p}_{ij}(u^{(i,1)}),\ldots,\mathfrak{p}_{ij}(u^{(i,r)})\}\,=\,\{\mathfrak{p}_{ji}(u^{(j,1)}),\ldots,\mathfrak{p}_{ji}(u^{(j,r)})\}.$
Up to permutation, we have
(3.14) $\begin{array}[]{c}\mathfrak{p}_{ij}(u^{(i,1)})=\mathfrak{p}_{ji}(u^{(j,1)}),\ldots,\mathfrak{p}_{ij}(u^{(i,r)})=\mathfrak{p}_{ji}(u^{(j,r)}),\\ \lambda_{i,1}=\lambda_{j,1},\ldots,\lambda_{i,r}=\lambda_{j,r}.\end{array}$
Since $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$, the above implies that $\lambda_{i,l}=\lambda_{j,l}$ for all $i,j,l$. Moreover, all $u^{(i,l)},u^{(j,l)}$ are projections of a common vector in $\mathbb{R}^{\Delta_{i}\cup\Delta_{j}}$. Therefore, there are points $u^{(1)},\ldots,u^{(r)}\in\mathbb{R}^{n}$ such that (3.13) holds. Since $(u^{(j)})_{\Delta_{i}}\in K_{\Delta_{i}}$ for all $i$, we have $u^{(j)}\in K$ for all $j$. So every $[u^{(j)}]_{spa,2t}$ is feasible for (3.3) with the relaxation order equal to $t$, and every $[u^{(j)}]_{spa,2k}$ is feasible for (3.3) with the relaxation order equal to $k$. Therefore,
$f_{k}^{smo}=\left<f,y^{*}|_{2t}\right>=\left<f,y^{*}|_{2k}\right>\leq\langle f,[u^{(j)}]_{spa,2k}\rangle=\langle f,[u^{(j)}]_{spa,2t}\rangle,$
for all $j=1,\ldots,r$. Moreover, the decomposition (3.13) implies
$f_{k}^{smo}=\left<f,y^{*}|_{2t}\right>=\lambda_{1}\langle f,[u^{(1)}]_{spa,2t}\rangle+\cdots+\lambda_{r}\langle f,[u^{(r)}]_{spa,2t}\rangle.$
Since each $\lambda_{i}>0$, the above implies that for all $j=1,\ldots,r$,
$f_{k}^{smo}=\left<f,y^{*}|_{2t}\right>=\langle f,[u^{(j)}]_{spa,2t}\rangle=f(u^{(j)})\geq f_{\min}.$
Since $f_{k}^{smo}\leq f_{\min}$, all $u^{(1)},\ldots,u^{(r)}$ must be minimizers of (1.1). ∎
We remark that if the set $\mathcal{D}=\{\Delta_{1},\ldots,\Delta_{m}\}$ is not a connected cover of $[n]$, then $\mathcal{D}$ can be expressed as a union of disjoint subsets, say, $\mathcal{D}=\mathcal{D}_{1}\cup\cdots\cup\mathcal{D}_{\ell},$ such that each $\mathcal{D}_{i}$ is a connected cover for a subset of $[n]$. Then, the sparse optimization problem (1.1) can be split into a union of smaller sized problems that do not have mutually joint variables.
4. Some sufficient conditions for tightness
This section gives some sufficient conditions for the sparse Moment-SOS hierarchy to be tight. The sparse hierarchy of (3.2)-(3.3) is tight if and only if (3.5) or (3.6) holds for some $k$. This leads to the following assumption.
Assumption 4.1.
There exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ satisfying
(4.1) $\boxed{\begin{array}[]{c}\ p_{1}+\cdots+p_{m}+f_{\min}=0,\\ f_{i}+p_{i}\geq 0~{}\text{on}~{}K_{\Delta_{i}},\,i=1,\ldots,m.\end{array}}$
We remark that (4.1) does not imply $f_{i}+p_{i}\in\mbox{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}]$. This section explores various conditions for Assumption 4.1 to hold and for the sparse Moment-SOS hierarchy of (3.2)-(3.3) to be tight.
4.1. The convex case
We consider the case that (1.1) is a convex optimization problem. For each $i=1,\ldots,m$, we write that
(4.2) $h_{i}=(h_{i,1},\ldots,h_{i,\ell_{i}}),\quad g_{i}=(g_{i,1},\ldots,g_{i,s_{i}}).$
Assume each $h_{i}$ is linear (or it does not appear) and $u$ is a minimizer of (1.1). Under certain constraint qualification like Slater’s condition (i.e., there is a feasible point $x$ for (1.1) such that all $g_{i}(x)>0$), there exist Lagrange multipliers $\lambda_{i,j}$ and $\nu_{i,j}$ such that ($\perp$ means that the product is 00)
(4.3) $\left\{\begin{gathered}\displaystyle\sum_{i=1}^{m}\nabla f_{i}(u_{\Delta_{i}})=\sum_{i=1}^{m}\Big{[}\sum_{j=1}^{s_{i}}\lambda_{i,j}\nabla g_{i,j}(u_{\Delta_{i}})+\sum_{j=1}^{\ell_{i}}\nu_{i,j}\nabla h_{i,j}(u_{\Delta_{i}})\Big{]},\\ 0\leq\lambda_{i,j}\perp g_{i,j}(u_{\Delta_{i}})\geq 0,\ i=1,\ldots,m,\ j=1,\ldots,s_{i}.\end{gathered}\right.$
This is called the first order optimality condition (FOOC).
Theorem 4.2.
Suppose that each $h_{i}$ is linear (or it does not appear), all $f_{i}$ and $-g_{ij}$ are convex, and $u$ is a minimizer of (1.1). If (4.3) holds, then Assumption 4.1 holds. Furthermore, if in addition, the sub-Hessian
$\nabla_{x_{\Delta_{i}}}^{2}\bigg{[}f_{i}(u_{\Delta_{i}})-\sum_{j=1}^{s_{i}}\lambda_{i,j}g_{i,j}(u_{\Delta_{i}})\bigg{]}$
is positive definite and each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}]$ is archimedean, then the sparse Moment-SOS hierarchy of (3.2)-(3.3) is tight.
Proof.
For each $i=1,\ldots,m$, let
$p_{i}(x)\coloneqq-(x-u)^{T}\big{[}\nabla f_{i}(u_{\Delta_{i}})-\sum_{j=1}^{s_{i}}\lambda_{i,j}\nabla g_{i}(u_{\Delta_{i}})-\sum_{j=1}^{\ell_{i}}\nu_{i,j}\nabla h_{i}(u_{\Delta_{i}})\big{]}-f_{i}(u_{\Delta_{i}}).$
Since $f_{i}$, $g_{i}$, $h_{i}$ only depend on $x_{\Delta_{i}}$, we have $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$. Note
$f_{1}(u_{\Delta_{1}})+\cdots+f_{m}(u_{\Delta_{m}})=f_{\min}.$
The equation in (4.3) and the above choice of $p_{i}$ imply
$p_{1}+\cdots+p_{m}+f_{\min}=0,$
$\nabla f_{i}(u_{\Delta_{i}})+\nabla p_{i}(u_{\Delta_{i}})=\sum_{j=1}^{s_{i}}\lambda_{i,j}\nabla g_{i}(u_{\Delta_{i}})+\sum_{j=1}^{\ell_{i}}\nu_{i,j}\nabla h_{i}(u_{\Delta_{i}}),$
so $u_{\Delta_{i}}$ satisfies the FOOC for
(4.4) $\left\{\begin{array}[]{cl}\displaystyle\min_{x_{\Delta_{i}}}&f_{i}(x_{\Delta_{i}})+p_{i}(x_{\Delta_{i}})\\ \mbox{s.t.}&h_{i}(x_{\Delta_{i}})=0,\,g_{i}(x_{\Delta_{i}})\geq 0.\end{array}\right.$
Since $p_{i}(x_{\Delta_{i}})$ is linear, $f_{i}(x_{\Delta_{i}})+p_{i}(x_{\Delta_{i}})$ is also convex in $x_{\Delta_{i}}$, hence $u_{\Delta_{i}}$ is a minimizer of (4.4). Its minimum value is 00, so $f_{i}+p_{i}\geq 0$ on $K_{\Delta_{i}}$. Therefore, Assumption 4.1 holds. The tightness of the hierarchy of (3.2)-(3.3) follows from Theorem 3.1 and [4, Corollary 3.3] (or [30, Theorem 7.2.5]). ∎
Theorems 3.3 and 3.4 can be applied to get minimizers of (1.1). Note that a convex optimization problem typically has a unique optimizer, or otherwise it has infinitely many optimizers. For this reason, when the conditions (3.8)-(3.9) and/or (3.12)-(3.13) hold, we typically have $r_{i}=1$ for all $i$. For such cases, if $y^{*}$ is a minimizer of (3.3), the point
$x^{*}\,\coloneqq\,(y^{*}_{e_{1}},\ldots,y^{*}_{e_{n}})$
is a minimizer of (1.1).
Next, we consider the special case that the defining polynomials of (1.1) are SOS-convex. Recall that a polynomial $\phi(x)$ is SOS-convex if its Hessian $\nabla^{2}\phi(x)=H(x)^{T}H(x)$ for some matrix polynomial $H(x)$. We refer to [30, Chap. 7] for SOS-convex polynomials. Under the SOS-convexity assumption, the sparse Moment-SOS relaxations (3.2)-(3.3) are tight for all $k\geq k_{0}$ and the condition (3.5) holds.
Theorem 4.3.
Assume the feasible set $K\neq\emptyset$ and the minimum value $f_{\min}>-\infty$. Suppose each $h_{i}$ is linear or it does not appear, each $f_{i}$ and $-g_{ij}$ are SOS-convex in $x_{\Delta_{i}}$. Then, we have:
  1. (i)
    For all $k\geq k_{0}$, $f^{smo}_{k}=f_{\min}$. Moreover, if Slater’s condition holds, then $f^{spa}_{k}=f_{\min}$ and
    (4.5) $f-f_{\min}\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}.$
  2. (ii)
    For every minimizer $y^{*}$ of (3.3), the point $x^{*}=(y^{*}_{e_{1}},\ldots,y^{*}_{e_{n}})$ is a minimizer of (1.1).
Proof.
Suppose $y$ is feasible for the relaxation (3.3) and let $v=(y_{e_{1}},\ldots,y_{e_{n}})$. Since they are SOS-convex, by Jensen’s inequality (see [16] or [30, Chap. 7]), we have
$f_{i}(v_{\Delta_{i}})\leq\langle f_{i},y_{\Delta_{i}}\rangle,\quad-g_{i,j}(v_{\Delta_{i}})\leq-\langle g_{i,j},y_{\Delta_{i}}\rangle,\quad i=1,\ldots,m.$
When $h_{i}$ is linear, $h_{i}(v_{\Delta_{i}})=\langle h_{i},y_{\Delta_{i}}\rangle$. The feasibility constraint in (3.3) implies that $g_{i,j}(v_{\Delta_{i}})\geq 0$, so $v$ is a feasible point for (1.1). Also note that
$f(v)=\sum_{i=1}^{m}f_{i}(v_{\Delta_{i}})\leq\sum_{i=1}^{m}\langle f_{i},y_{\Delta_{i}}\rangle=\langle f,y\rangle.$
The above holds for all $y$ that is feasible for (3.3), so $f_{\min}\leq f^{smo}_{k}$. On the other hand, we always have $f_{\min}\geq f^{smo}_{k}$, so $f^{smo}_{k}=f_{\min}$. Therefore, if $y^{*}$ is a minimizer of (3.3), then $x^{*}$ is a minimizer of (1.1).
When the Slater’s condition holds, the moment relaxation (3.3) has strictly feasible points. So, the strong duality holds between (3.2) and (3.3) and (3.2) achieves its optimal value. Therefore, $f^{spa}_{k}=f^{smo}_{k}=f_{\min}$ and (4.5) holds. ∎
4.2. The case of sufficient optimality conditions
Under the archimedianness, the dense Moment-SOS hierarchy in [14] is tight when linear independence constraint qualification (LICQC), strict complementarity condition (SCC) and second order sufficient condition (SOSC) hold at every minimizer, as shown in [29]. We refer to [30, Chapter 5.1] for these conditions. For sparse Moment-SOS relaxations, we have a similar conclusion under Assumption 4.1.
Theorem 4.4.
Suppose Assumption 4.1 holds and each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}]$ is archimedean in $\mathbb{R}[x_{\Delta_{i}}]$. Assume the LICQC, SCC and SOSC hold at every minimizer of the optimization problem
(4.6) $\left\{\begin{array}[]{rl}\min&f_{i}(x_{\Delta_{i}})+p_{i}(x_{\Delta_{i}})\\ \mbox{s.t.}&h_{i}(x_{\Delta_{i}})=0,\,g_{i}(x_{\Delta_{i}})\geq 0.\end{array}\right.$
Then, it holds that
  1. (i)
    When $k$ is big enough, we have $f^{spa}_{k}=f^{smo}_{k}=f_{\min}$. Moreover, if each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]$ is real radical, then
    (4.7) $f-f_{\min}\in\operatorname{Ideal}[h]_{spa}+\operatorname{QM}[g]_{spa}.$
  2. (ii)
    When $k$ is big enough, every minimizer $y^{*}$ of (3.3) satisfies the flat truncation condition: there exists a degree $t\in[k_{0},k]$ such that for every $i$,
    (4.8) $r_{i}=\mbox{rank}\,M_{\Delta_{i}}^{(t)}[y^{*}_{\Delta_{i}}]=\mbox{rank}\,M_{\Delta_{i}}^{(t-d_{i})}[y^{*}_{\Delta_{i}}],$
    where $d_{i}$ is as in (3.7). Therefore, the decomposition (3.8) holds and each point $x^{*}$ as in Theorem 3.3 is a minimizer of (1.1).
  3. (iii)
    Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$ (see the definition before Theorem 3.4). Assume (4.8) holds for all $i$ and let $\mathcal{X}_{\Delta_{i}}$ be the set as in (3.9). For all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, suppose the projection
    $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})=\{\mathfrak{p}_{ij}(u^{(i,1)}),\ldots,\mathfrak{p}_{ij}(u^{(i,r_{i})})\}$
    is a set of $r_{i}$ distinct points. Then, when $t$ is big enough, we have for all $i,j$
    (4.9) $r\coloneqq r_{1}=\cdots=r_{m}=\mbox{rank}\,M_{\Delta_{ij}}^{(t)}[y^{*}_{\Delta_{ij}}]=\mbox{rank}\,M_{\Delta_{ij}}^{(t-1)}[y^{*}_{\Delta_{ij}}],$
    and there exist $u^{(1)},\ldots,u^{(r)}\in K$ such that (see (3.11) for the notation)
    (4.10) $\boxed{\begin{gathered}y^{*}|_{spa,2t}=\lambda_{1}[u^{(1)}]_{spa,2t}+\cdots+\lambda_{r}[u^{(r)}]_{spa,2t},\\ \lambda_{1}>0,\ldots,\lambda_{r}>0,\lambda_{1}+\cdots+\lambda_{r}=1.\end{gathered}}$
    Moreover, $u^{(1)},\ldots,u^{(r)}$ are minimizers of (1.1).
Proof.
(i) Since each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}]$ is archimedean, the feasible sets $K_{\Delta_{i}}$ and $K$ are all compact. So, (1.1) achieves its minimum value, say, at a minimizer $u\in K$. Assumption 4.1 implies
$(f_{1}+p_{1})+\cdots+(f_{m}+p_{m})=f-f_{\min}.$
Since $f(u)=f_{\min}$ and each $f_{i}+p_{i}\geq 0$ on $K_{\Delta_{i}}$, the minimum value of (4.6) is 00. By [29, Theorem 1.1], if the LICQC, SCC, and SOSC hold at every minimizer of (4.6), then there is a relaxation order $N_{i}$ such that for all $\epsilon>0$,
(4.11) $f_{i}+p_{i}-(-\epsilon)\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2N_{i}}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2N_{i}}.$
Let $N\coloneqq\max\{N_{1},\ldots,N_{m}\}$. The above then implies
$f+p_{1}+\cdots+p_{m}+m\epsilon\in\operatorname{Ideal}[h]_{spa,2N}+\operatorname{QM}[g]_{spa,2N}.$
By Assumption 4.1, we have $p_{1}+\dots+p_{m}+f_{\min}=0$, so for all $k\geq N$
$f-(f_{\min}-m\epsilon)$ $=$ $f+\sum_{i=1}^{m}p_{i}+m\epsilon-\Big{(}\sum_{i=1}^{m}p_{i}+f_{\min}\Big{)}$
$\in$ $\operatorname{Ideal}[h]_{spa,k}+\operatorname{QM}[g]_{spa,k}.$
Thus, for arbitrary $\epsilon>0$, we have $f_{k}^{spa}\geq f_{\min}-m\epsilon.$ On the other hand, $f_{k}^{spa}\leq f_{\min}$, so this forces $f_{k}^{spa}=f_{\min}$.
Moreover, if each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]$ is real radical, then
(4.12) $f_{i}+p_{i}-0\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}].$
This follows from [23, Theorem 9.5.3] and [29, Theorem 3.1]. So, we get
$f+p_{1}+\ldots+p_{m}\in\operatorname{Ideal}[h]_{spa}+\operatorname{QM}[g]_{spa}.$
By Assumption 4.1, $p_{1}+\cdots+p_{m}+f_{\min}=0$, so
$f-f_{\min}=f+\sum_{i=1}^{m}p_{i}-\Big{(}\sum_{i=1}^{m}p_{i}+f_{\min}\Big{)}\in\operatorname{Ideal}[h]_{spa}+\operatorname{QM}[g]_{spa}.$
(ii) For each degree $k\geq k_{0}$, the $k$th order SOS relaxation for (4.6) is
(4.13) $\left\{\begin{array}[]{cll}\max&\gamma\\ \mbox{s.t.}&f_{i}+p_{i}-\gamma\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2k}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2k}.\end{array}\right.$
Its dual optimization is the moment relaxation
(4.14) $\left\{\begin{array}[]{cll}\min&\langle f_{i}+p_{i},y_{\Delta_{i}}\rangle\\ \mbox{s.t.}&\mathscr{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}]=0,L_{g_{i}}^{\Delta_{i},k}[y_{\Delta_{i}}]\succeq 0,\\ &(y_{\Delta_{i}})_{0}=1,\ M_{\Delta_{i}}^{(k)}[y_{\Delta_{i}}]\succeq 0,\\ &y_{\Delta_{i}}\in\mathbb{R}^{\mathbb{N}^{\Delta_{i}}_{2k}}.\end{array}\right.$
The minimum value of (4.6) is 00. Since each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}]$ is archimedean and the LICQC, SCC, SOSC hold at each optimizer of (4.6), the Moment-SOS hierarchy of (4.13)-(4.14) has finite convergence. This is shown in [29]. So, there exists $N_{0}$ such that for all $k\geq N_{0}$ and all $\epsilon>0$,
$f_{i}+p_{i}-(-\epsilon)\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2k}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2k}.$
Since $y^{*}$ is a minimizer of (3.3), $y^{*}_{\Delta_{i}}$ is feasible for (4.14). Hence,
$\langle f_{i}+p_{i},y^{*}_{\Delta_{i}}\rangle+\epsilon=\langle f_{i}+p_{i}+\epsilon,y^{*}_{\Delta_{i}}\rangle\geq 0.$
Since $\epsilon>0$ can be arbitrary, we get $\langle f_{i}+p_{i},y^{*}_{\Delta_{i}}\rangle\geq 0$. By item (i), it holds
00 0 $=$ $f^{smo}_{k}-f_{\min}=\langle f,y^{*}\rangle-f_{\min}=\langle f-f_{\min},y^{*}\rangle$
$=$ $\Big{\langle}\sum_{i=1}^{m}f_{i}+p_{i},y^{*}\Big{\rangle}=\sum_{i=1}^{m}\langle f_{i}+p_{i},y^{*}_{\Delta_{i}}\rangle.$
In the above, we have used $-f_{\min}=p_{1}+\cdots+p_{m}$ by Assumption 4.1. So, $\langle f_{i}+p_{i},y^{*}_{\Delta_{i}}\rangle=0$ for every $i$. This means that $y^{*}_{\Delta_{i}}$ is a minimizer for (4.14). Therefore, the conclusion follows from Theorem 3.3 of [9] and Theorem 3.3.
(iii) Since LICQC, SCC and SOSC hold at each minimizer of (4.6), the set $S_{i}$ of all minimizers of (4.6) is finite. Note that $\mathcal{X}_{\Delta_{i}}\subseteq S_{i}$, so $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})$ is also a finite set. Pick arbitrary $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, then $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})$ is a set of $r_{i}$ distinct points, by the given assumption. We know that $y^{*}_{\Delta_{ij}}$ has a representing measure whose support is $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})$. When $t$ is big enough, we must have
$\mbox{rank}\,M_{\Delta_{ij}}^{(t)}[y^{*}_{\Delta_{ij}}]=\mbox{rank}\,M_{\Delta_{ij}}^{(t-1)}[y^{*}_{\Delta_{ij}}].$
This is because $\mbox{rank}\,M_{\Delta_{ij}}^{(t)}[y^{*}_{\Delta_{ij}}]$ is uniformly bounded above by the cardinality of $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})$. So, the truncation $y^{*}_{\Delta_{ij}}|_{2t}$ is flat and it has a unique representing measure, say, $\mu_{ij}$. Since $y^{*}_{\Delta_{ij}}|_{2t}=y^{*}_{\Delta_{ji}}|_{2t}$, we have $\mu_{ij}=\mu_{ji}$. Hence, $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})=\mathfrak{p}_{ji}(\mathcal{X}_{\Delta_{j}})$. Since $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})$ consists of $r_{i}$ distinct points, we must have $r_{i}=r_{j}$. Therefore, $r_{1}=\cdots=r_{m}=r$, since $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$. Hence, (4.9) holds. The remaining conclusions follow from Theorem 3.4. ∎
4.3. The case of finite sets
We prove the sparse Moment-SOS hierarchy of (3.2)-(3.3) is tight when the equality constraint of (4.6) defines a finite real variety.
Theorem 4.5.
Suppose Assumption 4.1 holds and each real variety $V_{\mathbb{R}}(h_{i})$ is finite. Then, it holds that
  1. (i)
    When $k$ is big enough, we have $f^{spa}_{k}=f^{smo}_{k}=f_{\min}$. Moreover, if each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]$ is real radical, then
    (4.15) $f-f_{\min}\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}.$
  2. (ii)
    When $k$ is big enough, every minimizer $y^{*}$ of (3.3) satisfies the flat truncation condition: there exists a degree $t\in[k_{0},k]$ such that for every $i$,
    (4.16) $r_{i}=\mbox{rank}\,M_{\Delta_{i}}^{(t)}[y^{*}_{\Delta_{i}}]=\mbox{rank}\,M_{\Delta_{i}}^{(t-d_{i})}[y^{*}_{\Delta_{i}}],$
    where $d_{i}$ is as in (3.7). Therefore, the decomposition (3.8) holds and each point $x^{*}$ as in Theorem 3.3 is a minimizer of (1.1).
  3. (iii)
    Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$ (see the definition before Theorem 3.4). Assume (4.16) holds for all $i$ and let $\mathcal{X}_{\Delta_{i}}$ be the set as in (3.9). For all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, suppose the projection
    $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})=\{\mathfrak{p}_{ij}(u^{(i,1)}),\ldots,\mathfrak{p}_{ij}(u^{(i,r_{i})})\}$
    is a set of $r_{i}$ distinct points. Then, when $t$ is big enough, we have for all $i,j$,
    (4.17) $\begin{array}[]{c}r\coloneqq r_{1}=\cdots=r_{m}=\mbox{rank}\,M_{\Delta_{ij}}^{(t)}[y^{*}_{\Delta_{ij}}]=\mbox{rank}\,M_{\Delta_{ij}}^{(t-1)}[y^{*}_{\Delta_{ij}}],\end{array}$
    and there exist $u^{(1)},\ldots,u^{(r)}\in K$ such that (see (3.11) for the notation)
    (4.18) $\boxed{\begin{gathered}y^{*}|_{spa,2t}=\lambda_{1}[u^{(1)}]_{spa,2t}+\cdots+\lambda_{r}[u^{(r)}]_{spa,2t},\\ \lambda_{1}>0,\ldots,\lambda_{r}>0,\lambda_{1}+\cdots+\lambda_{r}=1.\end{gathered}}$
    Moreover, $u^{(1)},\ldots,u^{(r)}$ are minimizers of (1.1).
Proof.
(i) Since each $V_{\mathbb{R}}(h_{i})$ is finite, there exists $N_{i}$ such that for all $\epsilon>0$,
$f_{i}+p_{i}+\epsilon\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2N_{i}}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2N_{i}}.$
This is shown in the proof of Theorem 1.1 of [28]. Let $N\coloneqq\max\{N_{1},\ldots,N_{m}\}.$ Then, for all $\epsilon>0$, we have
$f-(f_{\min}-m\epsilon)$ $=$ $\sum_{i=1}^{m}\big{[}f_{i}+p_{i}+\epsilon\big{]}-\Big{(}\sum_{i=1}^{m}p_{i}+f_{\min}\Big{)}$
$\in$ $\operatorname{Ideal}[h]_{spa,2N}+\operatorname{QM}[g]_{spa,2N}.$
So, $\gamma=f_{\min}-m\epsilon$ is feasible for (3.2) with $k=N$. Hence, for all $\epsilon>0$,
$f_{\min}-m\epsilon\leq f_{N}^{spa}\leq f_{\min}.$
This forces $f_{k}^{spa}=f_{\min}$ for all $k\geq N$.
Furthermore, if each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]$ is real radical, then
$f_{i}+p_{i}\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2N_{i}}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2N_{i}},$
when $N_{i}$ is big enough. This is implied by Proposition 5.6.4 of [30]. So, (4.15) follows.
(ii) Consider the relaxations (4.13)-(4.14). Since $y^{*}$ is a minimizer of (3.3), $y^{*}_{\Delta_{i}}$ is a minimizer of (4.14). This can be similarly shown as for Theorem 4.4(ii). Since $V_{\mathbb{R}}(h_{i})$ is finite, (4.16) must be satisfied, by Theorem 5.6.1 of [30].
(iii) Note that $\mathcal{X}_{\Delta_{i}}\subseteq V_{\mathbb{R}}(h_{i})$, so $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})$ is also a finite set. The remaining proof is the same as for Theorem 4.4(iii). ∎
For each $t=2,\ldots,m$, denote the sets
(4.19) $\widehat{\Delta}_{t-1}$$\coloneqq$ $\Delta_{1}\cup\cdots\cup\Delta_{t-1},\quad I_{t}\coloneqq\widehat{\Delta}_{t-1}\cap\Delta_{t},$
(4.20) $K_{\widehat{\Delta}_{t-1}}$$\coloneqq$ $\{x_{\widehat{\Delta}_{t-1}}:x_{\Delta_{i}}\in K_{\Delta_{i}},\enspace i=1,\ldots,t-1\}.$
For a set $S\subset\mathbb{R}^{\Gamma}$ with $\Gamma\supseteq I$, we denote the projection
$S_{I}\,\coloneqq\,\{x_{I}:x_{\Gamma}\in S\}.$
Theorem 4.6.
Assume that $K\neq\emptyset$, all $K_{\Delta_{i}}$ are compact, and $\Delta_{1},\ldots,\Delta_{m}$ satisfy the RIP. If the projections $(K_{\widehat{\Delta}_{t-1}})_{I_{t}}$ and $(K_{\Delta_{t}})_{I_{t}}$ are both finite sets for all $2\leq t\leq m$, then Assumption 4.1 holds.
The proof of Theorem 4.6 is given in Appendix A. The following result is straightforward by Theorems 4.5 and 4.6.
Theorem 4.7.
Suppose the real variety $V_{\mathbb{R}}(h_{i})$ is a finite set for every $i\in[m]$ and $\Delta_{1},\ldots,\Delta_{m}$ satisfy the RIP. Then, Assumption 4.1 holds and all conclusions of Theorem 4.5 hold.
5. The Schmüdgen type sparse relaxations
We write the tuples $g_{i},h_{i}$ as in (4.2). The Schmüdgen type sparse SOS relaxation for solving (1.1) is
(5.1) $\left\{\begin{array}[]{rcl}f_{k}^{smg}\coloneqq&\max&\gamma\\ &\mbox{s.t.}&f-\gamma\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{Pre}[g]_{spa,2k}.\end{array}\right.$
The dual of (5.1) is the Schmüdgen type sparse moment relaxation
(5.2) $\left\{\begin{array}[]{ccl}f_{k}^{smm}\coloneqq&\min&\langle f,y\rangle=\langle f_{1},y_{\Delta_{1}}\rangle+\cdots+\langle f_{m},y_{\Delta_{m}}\rangle\\ &\mbox{s.t.}&\mathscr{V}_{h_{i}}^{spa,2k}[y]=0\ (i\in[m]),\\ &&L_{g_{i,J}}^{spa,k}[y]\succeq 0\ (i\in[m],\,J\subseteq[s_{i}]),\\ &&y_{0}=1,\ y\in\mathbb{R}^{\mathbb{U}_{k}}.\end{array}\right.$
We refer to (2.2) and (2.14) for the above notation. For a subset $J\subseteq[s_{i}]$, denote $g_{i,J}\coloneqq\prod_{j\in J}g_{i,j}.$ For the case $J=\emptyset$, $g_{i,\emptyset}=1$ and $L_{g_{i,\emptyset}}^{spa,k}[y]$ becomes the sparse moment matrix $M_{\Delta_{i}}^{(k)}[y_{\Delta_{i}}]$. In the same way as for Theorem 3.1, we can get the following by Lemma 2.1.
Theorem 5.1.
For the Schmüdgen type relaxations (5.1)-(5.2), we have:
  1. (i)
    For each $k\geq k_{0}$, it holds
    (5.3) $f-f_{\min}\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{Pre}[g]_{spa,2k}$
    if and only if there exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ such that
    (5.4) $\boxed{\begin{array}[]{c}\ p_{1}+\cdots+p_{m}+f_{\min}=0,\\ f_{i}+p_{i}\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2k}+\operatorname{Pre}_{\Delta_{i}}[g_{i}]_{2k},\,i\in[m].\end{array}}$
  2. (ii)
    When (5.4) holds for some $k$, the minimum value $f_{\min}$ of (1.1) is achievable if and only if there is a common zero point in $K$ for all $f_{i}+p_{i}$.
Now we consider the case that $K_{\Delta_{i}}$ is a finite set for every $i\in[m]$.
Theorem 5.2.
Suppose Assumption 4.1 holds and each $K_{\Delta_{i}}$ is a finite set. Then, we have:
  1. (i)
    The Schmüdgen type sparse Moment-SOS hierarchy of (5.1)-(5.2) is tight, i.e., $f_{k}^{smg}=f_{\min}$ when $k$ is big enough.
  2. (ii)
    When $k$ is large enough, every minimizer $y^{*}$ of (5.2) satisfies the flat truncation: there exists a degree $t\in[k_{0},k]$ such that for every $i$,
    (5.5) $r_{i}=\mbox{rank}\,M_{\Delta_{i}}^{(t)}[y^{*}_{\Delta_{i}}]=\mbox{rank}\,M_{\Delta_{i}}^{(t-d_{i})}[y^{*}_{\Delta_{i}}],$
    where $d_{i}$ is as in (3.7). Therefore, the decomposition (3.8) holds and each point $x^{*}$ as in Theorem 3.3 is a minimizer of (1.1).
  3. (iii)
    Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$ (see the definition before Theorem 3.4). Assume (5.5) holds for all $i$. Let $\mathcal{X}_{\Delta_{i}}$ be the set as in (3.9). For all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, suppose the projection
    $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})=\{\mathfrak{p}_{ij}(u^{(i,1)}),\ldots,\mathfrak{p}_{ij}(u^{(i,r_{i})})\}$
    is a set of $r_{i}$ distinct points. Then, when $t$ is big enough, we have for all $i,j$,
    (5.6) $r\coloneqq r_{1}=\cdots=r_{m}=\mbox{rank}\,M_{\Delta_{ij}}^{(t)}[y^{*}_{\Delta_{ij}}]=\mbox{rank}\,M_{\Delta_{ij}}^{(t-1)}[y^{*}_{\Delta_{ij}}],$
    and there exist $u^{(1)},\ldots,u^{(r)}\in K$ such that (see (3.11) for the notation)
    (5.7) $\boxed{\begin{gathered}y^{*}|_{spa,2t}=\lambda_{1}[u^{(1)}]_{spa,2t}+\cdots+\lambda_{r}[u^{(r)}]_{spa,2t},\\ \lambda_{1}>0,\ldots,\lambda_{r}>0,\quad\lambda_{1}+\cdots+\lambda_{r}=1.\end{gathered}}$
    Moreover, $u^{(1)},\ldots,u^{(r)}$ are minimizers of (1.1).
Proof.
(i) As we have shown in Theorem 4.4(i), the minimum value of (4.6) is 00. Since $K_{\Delta_{i}}$ is finite, there exists a degree $N_{i}$ such that for all $\epsilon>0$,
$f_{i}+p_{i}+\epsilon\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2N_{i}}+\mbox{Pre}_{\Delta_{i}}[g_{i}]_{2N_{i}}.$
This is shown in the proof of Theorem 4.1 of [28]. Note that
$\sum_{i=1}^{m}\left(f_{i}+p_{i}+\epsilon\right)=f+m\epsilon-f_{\min}+\Big{(}\sum_{i=1}^{m}p_{i}+f_{\min}\Big{)}.$
Let $N\coloneqq\max\left\{N_{1},\ldots,N_{m}\right\}.$ Then, for all $\epsilon>0$, we have
$f-(f_{\min}-m\epsilon)$ $=$ $\sum_{i=1}^{m}\left(f_{i}+p_{i}+\epsilon\right)-\Big{(}\sum_{i=1}^{m}p_{i}+f_{\min}\Big{)}$
$\in$ $\operatorname{Ideal}[h]_{spa,2N}+\operatorname{Pre}[g]_{spa,2N}.$
This implies that $\gamma=f_{\min}-m\epsilon$ is feasible for (5.1) with the order $k=N$. Hence, for all $\epsilon>0$,
$f_{\min}-m\epsilon\leq f_{N}^{smg}\leq f_{\min}.$
This forces $f_{N}^{smg}=f_{\min}$, so $f_{k}^{smg}=f_{\min}$ for all $k\geq N$.
(ii) Suppose $y^{*}$ is a minimizer of (5.2). In the same way as for Theorem 4.4(ii), we can show that each $y^{*}_{\Delta_{i}}$ is a minimizer of the moment relaxation
(5.8) $\left\{\begin{array}[]{cll}\min&\langle f_{i}+p_{i},y_{\Delta_{i}}\rangle\\ \mbox{s.t.}&\mathscr{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}]=0\,(i\in[m]),\\ &L_{g_{i,J}}^{\Delta_{i},k}[y_{\Delta_{i}}]\succeq 0\,(i\in[m],\,J\subseteq[s_{i}]),\\ &(y_{\Delta_{i}})_{0}=1,\ y_{\Delta_{i}}\in\mathbb{R}^{\mathbb{N}^{\Delta_{i}}_{2k}}.\end{array}\right.$
Since each $K_{\Delta_{i}}$ is finite, by [30, Theorem 5.6.7], we know $y^{*}_{\Delta_{i}}$ satisfies the flat truncation (5.5) when $k$ is large enough. The remaining part follows from Theorem 3.3.
(iii) Note that $\mathcal{X}_{\Delta_{i}}\subseteq K_{\Delta_{i}}$, so $\mathfrak{p}_{ij}(\mathcal{X}_{\Delta_{i}})$ is also a finite set. Then, by the same proof as for Theorem 4.4(iii), we can get (5.6). ∎
We remark that when all $K_{\Delta_{i}}$ are finite sets and the RIP holds, then Schmüdgen type sparse Moment-SOS hierarchy of (5.1)-(5.2) is tight.
Theorem 5.3.
Suppose all $K_{\Delta_{i}}$ are finite sets. If $\Delta_{1},\ldots,\Delta_{m}$ satisfy the RIP, then Assumption 4.1 holds and all conclusions in Theorem 5.2 hold.
Proof.
This follows directly from Theorem 4.6 and Theorem 5.2. ∎
6. Some examples
This section provides numerical experiments for the sparse Moment-SOS hierarchy of (3.2)-(3.3). For Examples 6.1-6.3, we use Yalmip [20] to implement sparse Moment-SOS relaxations, and apply Gloptipoly 3 [7] to check flat truncation conditions and extract minimizers. For sparse convex optimization problems in Examples 6.4-6.6, we apply the software TSSOS [21, 35, 37]. All semidefinite programs are solved by the software Mosek [1]. The computation is implemented in Julia 1.10.3/MATLAB 2023b, in an Apple MacBook Pro Laptop in macOS 14.2.1 with 12$\times$×Apple M3 Pro CPU and RAM 18GB. For neatness, only four decimal digits are displayed for computational results.
Example 6.1.
Consider the sparse optimization problem
(6.1) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=\underbrace{x_{1}^{2}x_{2}(x_{1}^{2}+x_{2}-1)}_{f_{1}}+\underbrace{x_{2}^{2}x_{3}(x_{2}^{2}+x_{3}-1)}_{f_{2}}\\ \mbox{s.t.}&1-x_{1}^{2}-x_{2}^{2}\geq 0,\quad 1-x_{2}^{2}-x_{3}^{2}\geq 0,\end{array}\right.$
with $\Delta_{1}=\{1,2\}$ and $\Delta_{2}=\{2,3\}$. By solving (3.2)-(3.3) with $k=3$, we get $f^{spa}_{3}\approx 0.0666$. The condition (3.5) is satisfied for
$p_{1}(x_{\Delta_{1}})$$\approx$ $0.04149+0.0426x_{2}-0.1275x_{2}^{2}-0.1107x_{2}^{3}+0.2197x_{2}^{4}$
$+0.0669x_{2}^{5}-0.1037x_{2}^{6},$
and $p_{2}=-f_{\min}-p_{1}$. For $t=k=3$, (3.8) holds for both $\Delta_{1}$ and $\Delta_{2}$, and
$\mathcal{X}_{\Delta_{1}}=\{(-0.5100,0.4798),(0.5100,0.4798)\},\quad\mathcal{X}_{\Delta_{2}}=\{(0.4798,0.3849)\}.$
By Theorem 3.3, we get two minimizers: (±$(\pm 0.5100,0.4798,0.3849).$
Example 6.2.
Consider the sparse optimization problem
(6.2) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=\underbrace{x_{1}^{2}+4x_{1}x_{2}}_{f_{1}}+\underbrace{4x_{2}x_{3}-x_{3}^{2}}_{f_{2}}\\ \mbox{s.t.}&1-x_{1}^{2}-x_{2}^{2}\geq 0,\quad 1-x_{2}^{2}-x_{3}^{2}\geq 0,\end{array}\right.$
with $\Delta_{1}=\{1,2\}$ and $\Delta_{2}=\{2,3\}$. Then, (3.5) is satisfied with
$p_{1}(x_{\Delta_{1}})=1.0000x_{2}^{2}+1.0000,\quad p_{2}(x_{\Delta_{2}})=-1.0000x_{2}^{2}+3.0000.$
We solve (3.2)-(3.3) with the order $k=2$. For the minimizer $y^{*}$ of (3.3), the flat truncation condition (3.10) is satisfied for both $\Delta_{1}$ and $\Delta_{2}$ with $r_{1}=r_{2}=2$, and (3.12) is satisfied with
$\mbox{rank}\,M_{\Delta_{12}}^{(1)}[y^{*}_{\Delta_{12}}]=\mbox{rank}\,M_{\Delta_{12}}^{(2)}[y^{*}_{\Delta_{12}}]=2.$
By Theorem 3.4, we get $f_{\min}=-4.0000$ and two minimizers:
±$\pm(0.7071,-0.7071,0.7071).$
Example 6.3.
Consider the sparse optimization problem
(6.3) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=\underbrace{x_{1}x_{2}x_{3}-x_{1}x_{2}}_{f_{1}}+\underbrace{x_{2}x_{3}x_{4}-x_{3}x_{4}}_{f_{2}}\\ \mbox{s.t.}&x_{1}^{2}=x_{1},\,x_{2}^{2}=x_{2},\,x_{3}^{2}=x_{3},\,x_{4}^{2}=x_{4},\\ &x_{1}+x_{2}+x_{3}\geq 1,x_{2}+x_{3}+x_{4}\geq 1,\end{array}\right.$
with $\Delta_{1}=\{1,2,3\}$, $\Delta_{2}=\{2,3,4\}$, and
$\begin{gathered}h_{1}=(x_{1}^{2}-x_{1},x_{2}^{2}-x_{2},x_{3}^{2}-x_{3}),\quad g_{1}=(x_{1}+x_{2}+x_{3}-1),\\ h_{2}=(x_{2}^{2}-x_{2},x_{3}^{2}-x_{3},x_{4}^{2}-x_{4}),\quad g_{2}=(x_{2}+x_{3}+x_{4}-1).\\ \end{gathered}$
We solve the sparse relaxation (3.2)-(3.3) with $k=2$. For the minimizer $y^{*}$ of (3.3), the condition (3.8) holds with $t=k=2$ for both $\Delta_{1}$ and $\Delta_{2}$. By Theorem 3.3, we get $f_{\min}=f^{smo}_{2}=-1.0000$ and four minimizers:
$\begin{gathered}(0.0000,0.0000,1.0000,1.0000),\quad(1.0000,0.0000,1.0000,1.0000),\\ (1.0000,1.0000,0.0000,0.0000),\quad(1.0000,1.0000,0.0000,1.0000).\\ \end{gathered}$
Moreover, the condition (3.5) is satisfied for
$p_{1}(x_{\Delta_{1}})$$\approx$ $0.5001+0.4195x_{2}-0.2165x_{3}-0.1484x_{2}x_{3}+0.1559x_{2}^{2}$
$-0.4111x_{3}^{2}-0.8565x_{2}^{3}+0.0166x_{2}^{2}x_{3}+0.1820x_{2}x_{3}^{2}-0.7216x_{3}^{3}$
$+0.6349x_{2}^{2}x_{3}^{2}-0.4251x_{2}^{3}x_{3}-0.2599x_{2}x_{3}^{3}+0.7810x_{2}^{4}+0.8491x_{3}^{4},$
and $p_{2}=f_{\min}-p_{1}$.
Example 6.4.
Consider the convex optimization problem:
(6.4) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=f_{1}(x_{1},x_{2})+f_{2}(x_{2},x_{3})+f_{3}(x_{1},x_{3})\\ \mbox{s.t.}&1-x_{1}^{4}-x_{2}^{4}\geq 0,\ 1-x_{2}^{4}-x_{3}^{4}\geq 0,\ 1-x^{4}_{1}-x_{3}^{4}\geq 0.\end{array}\right.$
In the above, $f_{1}\,\coloneqq\,x_{1}^{6}+x_{2}^{6}+x_{1}^{3}x_{2}^{3}+x_{1}$,
$\begin{gathered}f_{2}\,\coloneqq\,x_{2}^{6}+x_{3}^{6}+x_{2}^{3}x_{3}^{3}-x_{2},\quad f_{3}\,\coloneqq\,x_{1}^{6}+x_{3}^{6}+x_{1}^{3}x_{3}^{3}+2x_{3}.\end{gathered}$
Each $f_{i}$ is SOS-convex. For instance, $f_{1}$ is SOS-convex since
$\nabla^{2}f_{1}(x_{1},x_{2})=\begin{bmatrix}30x_{1}^{4}&9x_{1}^{2}x_{2}^{2}\\ 9x_{1}^{2}x_{2}^{2}&30x_{2}^{4}\end{bmatrix}=\begin{bmatrix}x_{1}^{2}&0\\ 0&x_{2}^{2}\end{bmatrix}\begin{bmatrix}30&9\\ 9&30\end{bmatrix}\begin{bmatrix}x_{1}^{2}&0\\ 0&x_{2}^{2}\end{bmatrix}.$
By Theorem 4.3, the sparse relaxation (3.2)-(3.3) is tight for all $k\geq 3$. Solving it with $k=3$, we get $f^{spa}_{3}\approx-2.2561.$ By Theorem 4.3 (ii), we get $f_{\min}=f^{spa}_{3}$ and the minimizer $(-0.6036,0.6852,-0.7092).$
Example 6.5.
Consider the convex quadratic optimization problem:
(6.5) $\left\{\begin{array}[]{cll}\displaystyle\min_{x\in\mathbb{R}^{n}}&\displaystyle\sum_{i=1}^{m}\big{[}x_{\Delta_{i}}^{T}Q_{i}x_{\Delta_{i}}+b_{i}^{T}x_{\Delta_{i}}\big{]}\\ \mbox{s.t.}&1-x_{\Delta_{i}}^{T}B_{i}x_{\Delta_{i}}-c_{i}^{T}x_{\Delta_{i}}\geq 0,\ i=1,\ldots,m.\end{array}\right.$
We set $m=n$ and each block $\Delta_{i}$ consists of $w$ elements as
(6.6) $\Delta_{i}=\left\{\begin{array}[]{ll}\{i,\ldots,i+w-1\}&\mbox{for }i\leq n-w+1,\\ \{i,\ldots,n,1,\ldots,w-n+i-1\}&\mbox{for }i>n-w+1.\end{array}\right.$
The $b_{i}$ and $c_{i}$ are randomly generated vectors obeying Gaussian distribution. Each $Q_{i}=R_{i}^{T}R_{i}$ is randomly generated, where the entries of $R_{i}$ obey Gaussian distribution. The matrix $B_{i}$ is generated in the same way. By Theorem 4.3, the sparse relaxation (3.2)-(3.3) is tight for all $k\geq 1$. We report the computational time (in seconds) for solving (6.5) by the sparse relaxation (3.2)-(3.3) and dense relaxation in [14] with order $k=1$. In Table 1, we display the time by the sparse relaxation (3.2)-(3.3) on the left and the time by the dense relaxation on the right, inside each parenthesis. For instance, when $n=100$ and $w=5$, the sparse relaxation took $0.04$ second, while the dense relaxation took $2.24$ seconds. When $n=500$, the dense relaxations cannot be solved since the computer is out of memory (oom).
Table 1. The computational time (in seconds) for solving (6.5) by the sparse Moment-SOS relaxations (left) and the dense Moment-SOS relaxation (right).
$n=100$ $n=200$ $n=300$ $n=500$
$w=5$ (0.04, 2.24) (0.10, 82.01) (0.17, 2205.36) (0.35, oom)
$w=10$ (0.11, 2.58) (0.28, 87.02) (0.44, 2221.67) (0.88, oom)
$w=20$ (0.54, 2.76) (1.81, 87.37) (2.03, 2477.07) (3.72, oom)
Example 6.6.
Consider the sparse convex polynomial optimization
(6.7) $\left\{\begin{array}[]{cll}\displaystyle\min_{x\in\mathbb{R}^{n}}&\displaystyle\sum_{i=1}^{m}\Big{[}b_{i}^{T}x_{\Delta_{i}}+x_{\Delta_{i}}^{T}Q_{i}x_{\Delta_{i}}+\big{(}{x_{\Delta_{i}}^{[2]}}\big{)}^{T}D_{i}x_{\Delta_{i}}^{[2]}\Big{]}\\ \mbox{s.t.}&1-c_{i}^{T}x_{\Delta_{i}}-x_{\Delta_{i}}^{T}B_{i}x_{\Delta_{i}}-\big{(}{x_{\Delta_{i}}^{[2]}}\big{)}^{T}H_{i}x_{\Delta_{i}}^{[2]}\geq 0,\ i=1,\ldots,m.\end{array}\right.$
In the above, for $x_{\Delta_{i}}=\big{[}x_{j_{1}}\ x_{j_{2}}\ \ldots\ x_{j_{n_{i}}}\big{]}^{T}$, we denote
$x_{\Delta_{i}}^{[2]}\,\coloneqq\,\big{[}x^{2}_{j_{1}}\ x^{2}_{j_{2}}\ \ldots\ x^{2}_{j_{n_{i}}}\big{]}^{T}.$
We set $m=n$ and each block $\Delta_{i}$ is given as in (6.6). All $b_{i},c_{i},Q_{i}$ and $B_{i}$ are randomly generated in the same way as in Example 6.5. Each matrix $D_{i}=R_{i}^{T}R_{i}$ is randomly generated with the entries of $R_{i}$ obeying the uniform distribution on $[0,1]$, and each matrix $H_{i}$ is generated in the same way. For these choices, the objective is a sum of SOS-convex polynomials and the constraining polynomials are SOS-concave. This can be shown as in [30, Example 7.1.4]. By Theorem 4.3, the sparse relaxation (3.2)-(3.3) is tight for all $k\geq 2$. We report the computational time (in seconds) for solving (6.7) by the sparse relaxation (3.2)-(3.3) and dense relaxation in [14] with $k=2$. In Table 2, we display the computational time by the sparse relaxation (3.2)-(3.3) on the left, and the time by the dense relaxation on the right, inside each parenthesis. When $n$ is $50$ or $100$, the dense relaxations cannot be solved since the computer is out of memory (oom).
Table 2. Computational time (in seconds) for solving (6.7) by the sparse Moment-SOS relaxations (left) and by the dense Moment-SOS relaxation (right).
$n=20$ $n=30$ $n=50$ $n=100$
$w=5$ (0.10, 17.53) (0.15, $1928.50$) (0.26, oom) (0.67, oom)
$w=8$ (0.39, 17.76) (0.63, $1978.11$) (1.05, oom) (2.16, oom)
$w=10$ (1.09, 18.82) (1.74, $2134.91$) (3.36, oom) (6.96, oom)
We conclude this section with an example such that the sparse Moment-SOS hierarchy (3.2)-(3.3) is not tight, while the dense one is tight.
Example 6.7.
Consider the optimization problem
(6.8) $\left\{\begin{array}[]{cl}\displaystyle\min_{x\in\mathbb{R}^{3}}&f(x)=\underbrace{x_{1}^{2}+(x_{1}x_{2}-1)^{2}}_{f_{1}}+\underbrace{(x_{2}x_{3})^{2}+(x_{3}-1)^{2}}_{f_{2}}\\ \mbox{s.t.}&1-x_{1}^{2}\geq 0,\ 1-x_{2}^{2}\geq 0,\,1-x_{3}^{2}\geq 0.\end{array}\right.$
Here, $\Delta_{1}=\{1,2\}$ and $\Delta_{2}=\{2,3\}$. For fixed $x_{2}\in[-1,1]$, the minimum value of $f_{1}(x_{1},x_{2})$ over $x_{1}\in[-1,1]$ is $f_{1,\min}=(x_{2}^{2}+1)^{-1}$, which is attained at $x_{1}=x_{2}(1+x_{2}^{2})^{-1}$. Similarly, the minimum value of $f_{2}(x_{2},x_{3})$ over $x_{3}\in[-1,1]$ is $f_{2,\min}=x_{2}^{2}(x_{2}^{2}+1)^{-1},$ which is attained at $x_{3}=(1+x_{2}^{2})^{-1}$. Hence, we have
$f_{\min}=f_{1,\min}+f_{2,\min}=1.$
However, Assumption 4.1 fails to hold. Suppose otherwise it holds, then there exist $p_{1}\in\mathbb{R}[x_{1},x_{2}]$ and $p_{2}\in\mathbb{R}[x_{2},x_{3}]$ such that $p_{1}+p_{2}+1=0$ and
(6.9) $f_{1}+p_{1}\geq 0\ \mbox{on}\ [-1,1]^{2},\quad f_{2}+p_{2}\geq 0\ \mbox{on}\ [-1,1]^{2}.$
Since $x_{2}$ is the only joint variable of $x_{\Delta_{1}}$ and $x_{\Delta_{2}}$, both $p_{1}$ and $p_{2}$ depend only on $x_{2}$. So, by the first condition of (6.9), we have
(6.10) $f_{1,\min}+p_{1}(x_{2})=(x_{2}^{2}+1)^{-1}+p_{1}(x_{2})\geq 0\quad\forall x_{2}\in[-1,1].$
Since $p_{2}=-1-p_{1}$, the second condition of (6.9) implies
(6.11) $f_{2,\min}-1-p_{1}(x_{2})=-(x_{2}^{2}+1)^{-1}-p_{1}(x_{2})\geq 0\quad\forall x_{2}\in[-1,1].$
Combining (6.10) and (6.11), we get
$p_{1}(x_{2})=(x_{2}^{2}+1)^{-1}\quad\forall x_{2}\in[-1,1].$
However, the above cannot hold since $p_{1}$ is a polynomial. Therefore, Assumption 4.1 does not hold for (6.8), and hence the sparse Moment-SOS hierarchy of (3.2)-(3.3) is not tight. On the other hand, one can verify that
$f(x)-f_{\min}=(x_{1}x_{2}+x_{3}-1)^{2}+(x_{1}-x_{2}x_{3})^{2},$
which shows that the dense Moment-SOS relaxation is tight.
7. Conclusions and Discussions
This paper studies the sparse Moment-SOS hierarchy of relaxations (3.2)-(3.3) for solving sparse polynomial optimization problem (1.1). We show that this sparse Moment-SOS hierarchy is tight if and only if (3.5) or (3.6) holds, i.e., the objective can be equivalently written as a sum of sparse nonnegative polynomials, each of which belongs to the sum of the ideal and quadratic module generated by the corresponding constraints. Under Assumption 4.1, we give some sufficient conditions for this sparse hierarchy to be tight: optimality conditions for (4.6) or finiteness of individual constraining sets. We also prove some conditions for Assumption 4.1 to hold. In particular, we show that the sparse Moment-SOS hierarchy is tight under some convexity assumptions.
Here are some interesting questions for future work.
  • When the RIP holds, if $K$ is a finite set but each $K_{\Delta_{i}}$ may not, is the Schmüdgen-type sparse hierarchy of (5.1)-(5.2) always tight?
  • Does Assumption 4.1 hold when $f_{i}(x_{\Delta_{i}})$, $g_{i}(x_{\Delta_{i}})$, $h_{i}(x_{\Delta_{i}})$ are generic polynomials?
  • When the RIP holds, does the sparse Positivstellensatz hold? That is, when the RIP holds and $K=\emptyset$, do there exist $\sigma\in\operatorname{Pre}[g]_{spa}$ and $\phi\in\operatorname{Ideal}[h]_{spa}$ such that (2.6) holds?
Acknowledgements
This project was begun at a SQuaRE at the American Institute of Mathematics (AIM). The authors would like to thank AIM for providing a supportive and mathematically rich environment. Jiawang Nie is partially supported by the NSF grant DMS-2110780. Zheng Qu is partially supported by the Hong Kong Research Grants Council GRF No. 17317122. Xindong Tang is partially supported by the Hong Kong Research Grants Council HKBU-15303423. Linghao Zhang is partially supported by the UCSD School of Physical Sciences Undergraduate Summer Research Award.
Appendix A Proof of Theorem 4.6
We first prove Theorem 4.6 for the case $m=2$.
Lemma A.1.
Consider (1.1) with $m=2$. Assume $f_{\min}>-\infty$, its feasible set $K\neq\emptyset$ and the following two optimization problems are bounded below:
(A.1) $\left\{\begin{array}[]{cll}\displaystyle\min&f_{1}(x_{\Delta_{1}})\\ \mbox{s.t.}&h_{1}(x_{\Delta_{1}})=0,\\ &g_{1}(x_{\Delta_{1}})\geq 0,\end{array}\right.\quad\left\{\begin{array}[]{cll}\displaystyle\min&f_{2}(x_{\Delta_{2}})\\ \mbox{s.t.}&h_{2}(x_{\Delta_{2}})=0,\\ &g_{2}(x_{\Delta_{2}})\geq 0.\end{array}\right.$
If the projections $\mathfrak{p}_{12}(K_{\Delta_{1}})$ and $\mathfrak{p}_{21}(K_{\Delta_{2}})$ are both finite sets, then there exists $p\in\mathbb{R}[x_{\Delta_{12}}]$ such that
(A.2) $f_{1}+p\geq 0\,\mbox{ on }\,K_{\Delta_{1}},\quad f_{2}-p-f_{\min}\geq 0\,\mbox{ on }\,K_{\Delta_{2}},$
the minimal values of $f_{1}+p$ on $K_{\Delta_{1}}$ and $f_{2}-p-f_{\min}$ on $K_{\Delta_{2}}$ are both zeros.
Proof.
By the assumption, we can write that
$\mathfrak{p}_{12}(K_{\Delta_{1}})\cup\mathfrak{p}_{21}(K_{\Delta_{2}})=\{u^{(1)},\ldots,u^{(D)}\}\subseteq\mathbb{R}^{{\Delta_{12}}}.$
Let $\varphi_{1},\ldots,\varphi_{D}\in\mathbb{R}[x_{\Delta_{12}}]$ be interpolating polynomials such that $\varphi_{i}(u^{(j)})=0$ for $i\neq j$ and $\varphi_{i}(u^{(j)})=1$ for $i=j$. Denote the sets
(A.3) $U_{0}\coloneqq\mathfrak{p}_{12}(K_{\Delta_{1}})\cap\mathfrak{p}_{21}(K_{\Delta_{2}}),\enspace U_{1}\coloneqq\mathfrak{p}_{12}(K_{\Delta_{1}})\backslash U_{0},\enspace U_{2}\coloneqq\mathfrak{p}_{21}(K_{\Delta_{2}})\backslash U_{0}.$
Define the optimal value functions:
(A.6) $\left\{\begin{array}[]{rcl}F_{1}(u)\coloneqq&\displaystyle\min&f_{1}(x_{\Delta_{1}})\\ &\mbox{s.t.}&x_{\Delta_{1}}\in K_{\Delta_{1}},\,x_{\Delta_{12}}=u,\end{array}\right.$
(A.9) $\left\{\begin{array}[]{rcl}F_{2}(u)\coloneqq&\displaystyle\min&f_{2}(x_{\Delta_{2}})\\ &\mbox{s.t.}&x_{\Delta_{2}}\in K_{\Delta_{2}},\,x_{\Delta_{12}}=u.\end{array}\right.$
For each $u\in U_{0}\cup U_{1}$, the problem (A.6) is feasible and its optimal value is finite. Similarly, for each $u\in U_{0}\cup U_{2}$, the problem (A.9) is feasible and its optimal value is finite. For all $u\in U_{0}$, $F_{1}(u)+F_{2}(u)$ equals the minimum value of
$\left\{\begin{array}[]{cll}\displaystyle\min&f_{1}(x_{\Delta_{1}})+f_{2}(x_{\Delta_{2}})\\ \mbox{s.t.}&x_{\Delta_{1}}\in K_{\Delta_{1}},\,x_{\Delta_{2}}\in K_{\Delta_{2}},\,x_{\Delta_{12}}=u.\end{array}\right.$
Since (1.1) is feasible and $U_{0}$ is given as in (A.3), we have
(A.10) $\min_{u\in U_{0}}[F_{1}(u)+F_{2}(u)]\,\,=\,\,f_{\min}.$
For $i=1,\ldots,D$, denote the values
$v_{i}\coloneqq\left\{\begin{array}[]{rl}-F_{1}(u^{(i)})&\mathrm{~{}if~{}}u^{(i)}\in U_{0}\cup U_{1},\\ F_{2}(u^{(i)})-f_{\min}&\mathrm{~{}if~{}}u^{(i)}\in U_{2}.\end{array}\right.$
Let $p\coloneqq v_{1}\varphi_{1}(x_{\Delta_{12}})+\cdots+v_{D}\varphi_{D}(x_{\Delta_{12}}),$ then $p\in\mathbb{R}[x_{\Delta_{12}}]$ and
(A.11) $p(u)=\left\{\begin{array}[]{rl}-F_{1}(u)&\mathrm{~{}if~{}}u\in U_{0}\cup U_{1},\\ F_{2}(u)-f_{\min}&\mathrm{~{}if~{}}u\in U_{2}.\end{array}\right.$
Observe the relations
$\min_{x_{\Delta_{1}}\in K_{\Delta_{1}}}\big{[}f_{1}(x_{\Delta_{1}})+p(x_{\Delta_{12}})\big{]}=\min_{u\in\mathfrak{p}_{12}(K_{\Delta_{1}})}\big{[}F_{1}(u)+p(u)\big{]},$
$\min_{x_{\Delta_{2}}\in K_{\Delta_{2}}}\big{[}f_{2}(x_{\Delta_{2}})-p(x_{\Delta_{12}})-f_{\min}\big{]}=\min_{u\in\mathfrak{p}_{21}(K_{\Delta_{2}})}\big{[}F_{2}(u)-p(u)-f_{\min}\big{]}.$
Since $\mathfrak{p}_{12}(K_{\Delta_{1}})=U_{0}\cup U_{1}$, (A.11) implies
[…]
Since $\mathfrak{p}_{21}(K_{\Delta_{2}})=U_{0}\cup U_{2}$, it holds
$\min_{u\in\mathfrak{p}_{21}(K_{\Delta_{2}})}[F_{2}(u)-p(u)-f_{\min}]\,=\,\min_{u\in U_{0}\cup U_{2}}[F_{2}(u)-p(u)-f_{\min}].$
Also note that
$F_{2}(u)-p(u)=f_{\min}\,\,\mbox{if}\,\,u\in U_{2},\quad F_{2}(u)-p(u)=F_{1}(u)+F_{2}(u)\,\,\mbox{if}\,\,u\in U_{0}.$
Then (A.10) implies
$\min_{u\in U_{2}}[F_{2}(u)-p(u)-f_{\min}]\,=\,\min_{u\in U_{0}}[F_{1}(u)+F_{2}(u)-f_{\min}]=0.$
This completes the proof. ∎
In the following, we prove Theorem 4.6 by induction.
Proof of Theorem 4.6..
The conclusion holds when $m=2$, by Lemma A.1. Suppose it holds for $m=k$. We prove it also holds for $m=k+1$. The problem (1.1) can be viewed to have two sparsity blocks: $\widehat{\Delta}_{m-1}$ and $\Delta_{m}$ (see (4.19) for the notation). Consider the following two optimization problems:
(A.12) $\left\{\begin{array}[]{cl}\min&f_{1}(x_{\Delta_{1}})+\cdots+f_{m-1}(x_{\Delta_{m-1}})\\ \mbox{s.t.}&x_{\Delta_{i}}\in K_{\Delta_{i}},\,\,i=1,\ldots,m-1,\end{array}\right.\quad\left\{\begin{array}[]{cl}\min&f_{m}(x_{\Delta_{m}})\\ \mbox{s.t.}&x_{\Delta_{m}}\in K_{\Delta_{m}}.\end{array}\right.$
The feasible set for the above left optimization problem is $K_{\widehat{\Delta}_{m-1}}$ (see (4.20) for the notation). By the assumption, both $K_{\widehat{\Delta}_{m-1}}$ and $K_{\Delta_{m}}$ are nonempty and compact; the optimal values of both optimization problems in (A.12) are finite. By Lemma A.1, there exists $p\in\mathbb{R}[x_{\widehat{\Delta}_{m-1}}]\cap\mathbb{R}[x_{\Delta_{m}}]$ such that
(A.13) $\left\{\begin{array}[]{ll}f_{1}+\cdots+f_{m-1}+p\geq 0\quad\mbox{on}\quad K_{\widehat{\Delta}_{m-1}},\\ f_{m}-p-f_{\min}\geq 0\quad\mbox{on}\quad K_{\Delta_{m}}.\end{array}\right.$
Furthermore, the minimum values of $f_{1}+\cdots+f_{m-1}+p$ on $K_{\widehat{\Delta}_{m-1}}$ and $f_{m}-p-f_{\min}$ on $K_{\Delta_{m}}$ are both equal to 00. The RIP ensures that there exists $i^{*}\in[m-1]$ such that $\widehat{\Delta}_{m-1}\cap\Delta_{m}\subseteq\Delta_{i^{*}}$. For $i=1,\ldots,m-1$, let
$\tilde{f}_{i}\coloneqq\left\{\begin{array}[]{ll}f_{i}&\enspace\mathrm{~{}if~{}}i\neq i^{*},\\ f_{i}+p&\enspace\mathrm{~{}if~{}}i=i^{*}.\end{array}\right.$
Then $\tilde{f}_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ for all $i=1,\ldots,m-1$ and the minimum value of $\tilde{f}_{1}+\cdots+\tilde{f}_{m-1}$ on $K_{\widehat{\Delta}_{m-1}}$ is zero. By the induction hypothesis, there exist polynomials $\tilde{q}_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ such that
(A.14) $\left\{\begin{gathered}\tilde{q}_{1}+\cdots+\tilde{q}_{m-1}=0,\\ \tilde{f}_{i}+\tilde{q}_{i}\geq 0~{}\text{on}~{}K_{\Delta_{i}},\,i=1,\ldots,m-1.\end{gathered}\right.$
Let $q_{m}\coloneqq-p-f_{\min}$, and for each $i=1,\ldots,m-1$, let
$q_{i}\coloneqq\left\{\begin{array}[]{ll}\tilde{q}_{i},&\enspace\mathrm{~{}if~{}}i\neq i^{*},\\ \tilde{q}_{i}+p,&\enspace\mathrm{~{}if~{}}i=i^{*}.\end{array}\right.$
Then, we have $q_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ for every $i\in[m]$ and
$\left\{\begin{gathered}q_{1}+\cdots+q_{m}=\tilde{q}_{1}+\cdots+\tilde{q}_{m-1}-f_{\min},\\ f_{m}+q_{m}=f_{m}-p-f_{\min},\\ f_{i}+q_{i}=\tilde{f}_{i}+\tilde{q}_{i},\enspace i=1,\ldots,m-1.\end{gathered}\right.$
Therefore, (A.13) and (A.14) imply that
(A.15) $\left\{\begin{gathered}q_{1}+\cdots+q_{m}+f_{\min}=0,\\ f_{i}+q_{i}\geq 0~{}\text{on}~{}K_{\Delta_{i}},\,i=1,\ldots,m.\end{gathered}\right.$
This completes the proof. ∎